### v1

# Energy-Conserving Data-Driven Discretizations for Maxwell’s Equations

Abstract
We study data-driven construction of spatial discretizations for the one-dimensional Maxwell system. Given high-fidelity training data generated by a spectral discretization, we learn a linear convolution stencil that approximates the spatial derivative operator appearing in Maxwell’s equations. The stencil is obtained by solving a convex quadratic optimization problem, subject to linear constraints that enforce skew-adjointness of the discrete derivative. These constraints guarantee a semi-discrete energy identity for the resulting Maxwell system. We prove that our constraints characterize the class of skew-symmetric convolution operators and express the associated numerical wave speed and CFL restriction for the classical leapfrog scheme in terms of the learned stencil’s Fourier symbol. We then compare several convex solvers for the resulting quadratic program—projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point reference implemented in CVXPY—and evaluate the learned schemes in time-dependent one-dimensional Maxwell simulations using a Crank–Nicolson (CN) time discretization. Our numerical experiments show that (i) energy-constrained learned stencils achieve accuracy comparable to standard central differences while exactly preserving the discrete electromagnetic energy under CN time-stepping, and (ii) ADMM and interior-point methods produce nearly identical operators, with ADMM offering a favorable tradeoff between accuracy, constraint satisfaction, and runtime.
1 Introduction
Finite-difference time-domain (FDTD) methods are a standard tool for solving Maxwell’s equations in computational electromagnetics [taf-love]. Classical schemes, such as the Yee method, are derived by hand from the continuous equations, with spatial derivatives approximated using fixed central-difference stencils and stability enforced via analytical CFL conditions. Recently, there has been growing interest in data-driven discretizations, where discrete operators are learned from numerical or experimental data, e.g., via regression or neural networks. Examples include data-driven discretization of Burgers and Navier–Stokes equations and learned coarse-grid operators for PDEs [bar-sinai].
Most existing data-driven schemes aim to minimize approximation error but do not explicitly enforce structural properties such as stability, conservation, or energy balance. For Maxwell’s equations this is a significant limitation, since standard FDTD schemes are carefully designed so that discrete versions of divergence constraints and energy identities hold at the semi-discrete level.
In this work we address this gap in a simple but concrete setting: the one-dimensional TE Maxwell system with periodic boundary conditions. Our goal is to learn a discrete spatial derivative operator
$D:\mathbb{R}^{N}\to\mathbb{R}^{N}$
represented as a convolution stencil, directly from data. We impose linear constraints on the stencil that enforce skew-adjointness of $D$ with respect to the discrete $L^{2}$ inner product. This structural constraint guarantees exact conservation of discrete electromagnetic energy at the semi-discrete level, regardless of the details of the training data.
Novelty and contributions.
Building on this context, our main contributions are as follows:
  • We formulate a convex quadratic optimization problem for learning convolution stencils for the 1D Maxwell system from spectral training data. The objective is a least-squares misfit between the learned discrete derivative and high-fidelity derivatives, while the constraints encode skew-symmetry of the stencil coefficients together with simple box bounds.
  • We show that these linear constraints are exactly equivalent to skew-adjointness of the corresponding convolution operator with respect to the discrete $L^{2}$ inner product, and hence to a semi-discrete electromagnetic energy identity. This yields, to our knowledge, the first learned Maxwell stencil with a guaranteed semi-discrete energy conservation law.
  • We characterize the numerical wave speed and associated CFL restriction of the learned discrete curl operator purely in terms of its Fourier symbol $\mu(\theta)$.
  • We implement projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point method (via CVXPY) to solve the constrained quadratic program, and we benchmark the resulting stencils in time-dependent 1D Maxwell simulations. The experiments compare PDE error, energy-constraint residuals, and energy evolution across stencil radii, and demonstrate that ADMM and interior-point methods produce nearly identical, highly accurate stencils, with ADMM offering a favorable tradeoff between accuracy, constraint satisfaction, and runtime.
Although our experiments focus on a simple 1D periodic model, the framework and energy constraints extend naturally to higher-dimensional settings and more complex material models.
Our results illustrate how even simple linear constraints in a data-driven discretization can encode important physics and restore structural properties such as energy conservation.
1.1 Related work
Data-driven discretizations.
There is a growing literature on learning discretizations and coarse-grid operators for PDEs directly from data. A prominent example is the work of [bar-sinai], who learn finite-difference schemes for Burgers and Navier–Stokes equations by fitting local stencils to high-fidelity solutions while optimizing accuracy on a training set. Related approaches use regression or neural networks to construct reduced models or learned fluxes for time-dependent PDEs; see, e.g., [parish2016paradigm, ling2016reynolds, rudy2017data, raissi2018hidden, kochkov2021machine]. In most of these methods, the learned operators are chosen solely to minimize an empirical error measure and are not required to satisfy structural constraints such as conservation, stability, or symmetries of the underlying PDE. In contrast, our focus is on learning a discrete Maxwell operator that provably satisfies a semi-discrete energy identity through an explicit skew-adjointness constraint on the convolution stencil.
Energy-preserving FDTD and skew-adjoint schemes.
Classical finite-difference time-domain (FDTD) methods for Maxwell’s equations, such as the Yee scheme and its variants, are carefully constructed so that discrete versions of divergence constraints and energy identities hold at the semi-discrete or fully discrete level [taf-love, richtmyer1967, morton1994, leveque2007]. In particular, staggered spatial discretizations and symplectic or A-stable time integrators such as the Crank–Nicolson scheme are often used to obtain skew-adjoint or nearly skew-adjoint update operators, leading to strong stability and conservation properties [crank1947, giles2010, hairerwanner2010]. There is also a large body of work on structure-preserving and geometric integrators for Hamiltonian ODE/PDE systems, which emphasizes invariants such as energy and symplectic form [hairer]. Our work is closest in spirit to these structure-preserving schemes, but differs in that the discrete spatial operator is learned from data rather than hand-designed: we encode skew-adjointness as a set of linear constraints on the stencil coefficients and solve a convex quadratic program to obtain an energy-conserving discretization.
This work in context.
To the best of our knowledge, the present paper is the first to combine data-driven stencil learning for Maxwell’s equations with linear constraints that (i) exactly characterize skew-symmetric convolution operators and (ii) yield a rigorous semi-discrete energy identity and a CFL condition expressed directly in terms of the learned stencil’s Fourier symbol. This provides a bridge between unconstrained data-driven discretizations and classical energy-preserving FDTD schemes, while remaining compatible with standard convex optimization toolchains (projected gradient, Nesterov acceleration, ADMM, and interior-point solvers).
2 Semi-discrete Maxwell system and energy identity
We consider the 1D Maxwell system in vacuum, written in non-dimensional form with $\varepsilon_{0}=\mu_{0}=1$:
$\partial_{t}E=\partial_{x}H,\qquad\partial_{t}H=\partial_{x}E,$ (1)
on the periodic domain $[0,L]$.
We discretize space using a uniform grid $x_{i}=i\Delta x$, $i=0,\dots,N-1$, with $\Delta x=L/N$, and collect the nodal values of the electric and magnetic fields into vectors
$\mathbf{E}(t)=(E_{0}(t),\dots,E_{N-1}(t))^{\top},\qquad\mathbf{H}(t)=(H_{0}(t),\dots,H_{N-1}(t))^{\top}.$
Let $D:\mathbb{R}^{N}\to\mathbb{R}^{N}$ denote a linear discrete derivative operator. In this work, $D$ is a convolution operator defined by a stencil $w\in\mathbb{R}^{2R+1}$,
$(D\mathbf{u})_{i}=\sum_{\ell=-R}^{R}w_{\ell}\,u_{i+\ell},\qquad i=0,\dots,N-1,$ (2)
with periodic wrap-around in the index $i+\ell$. The semi-discrete Maxwell system then reads
$\frac{d}{dt}\mathbf{E}(t)=D\mathbf{H}(t),\qquad\frac{d}{dt}\mathbf{H}(t)=D\mathbf{E}(t).$ (3)
We equip $\mathbb{R}^{N}$ with the standard discrete $L^{2}$ inner product
$\langle\mathbf{u},\mathbf{v}\rangle:=\Delta x\sum_{i=0}^{N-1}u_{i}v_{i},\qquad\|\mathbf{u}\|^{2}:=\langle\mathbf{u},\mathbf{u}\rangle.$ (4)
The corresponding discrete electromagnetic energy is
$\mathcal{E}(t):=\frac{1}{2}\|\mathbf{E}(t)\|^{2}+\frac{1}{2}\|\mathbf{H}(t)\|^{2}.$ (5)
Theorem 1 (Semi-discrete energy identity).
Let $D$ be any linear operator on $\mathbb{R}^{N}$, and let $\mathbf{E}(t),\mathbf{H}(t)$ satisfy (3). Then
$\frac{d}{dt}\mathcal{E}(t)=\langle D\mathbf{H}(t),\mathbf{E}(t)\rangle+\langle D\mathbf{E}(t),\mathbf{H}(t)\rangle.$ (6)
In particular, if $D$ is skew-adjoint with respect to (4), i.e.,
$\langle D\mathbf{u},\mathbf{v}\rangle=-\langle\mathbf{u},D\mathbf{v}\rangle\quad\forall\,\mathbf{u},\mathbf{v},$ (7)
then the discrete energy is exactly conserved:
$\frac{d}{dt}\mathcal{E}(t)=0.$ (8)
Proof.
Differentiating (5) in time and using (3) gives
$\frac{d}{dt}\mathcal{E}(t)$ $=\left\langle\frac{d}{dt}\mathbf{E}(t),\mathbf{E}(t)\right\rangle+\left\langle\frac{d}{dt}\mathbf{H}(t),\mathbf{H}(t)\right\rangle$
$=\langle D\mathbf{H}(t),\mathbf{E}(t)\rangle+\langle D\mathbf{E}(t),\mathbf{H}(t)\rangle,$
which is (6). If $D$ is skew-adjoint, then
$\langle D\mathbf{H},\mathbf{E}\rangle=-\langle\mathbf{H},D\mathbf{E}\rangle=-\langle D\mathbf{E},\mathbf{H}\rangle,$
so the two terms in (6) cancel and $\frac{d}{dt}\mathcal{E}(t)=0$. ∎
2.1 Skew-symmetric convolution operators and stencil constraints
We now specialize to convolution operators of the form (2) with periodic boundary conditions. Let $w=(w_{-R},\dots,w_{-1},w_{0},w_{1},\dots,w_{R})^{\top}$ be the stencil.
Lemma 1.
Let $D$ be defined by (2) with periodic wrap-around. Then the matrix entries of $D$ satisfy
$D_{ij}=w_{i-j},\qquad i,j=0,\dots,N-1,$ (9)
where the difference iji-j is taken modulo $N$.
Proof.
By definition,
$(D\mathbf{u})_{i}=\sum_{\ell=-R}^{R}w_{\ell}\,u_{i+\ell}=\sum_{j=0}^{N-1}w_{i-j}\,u_{j},$
where $j=i+\ell$ (modulo $N$). Hence $D_{ij}=w_{i-j}$. ∎
Theorem 2 (Characterization of skew-symmetric stencils).
Let $D$ be the convolution operator (2) with periodic boundary conditions and stencil $w\in\mathbb{R}^{2R+1}$. Then $D$ is skew-adjoint with respect to (4), i.e., $D^{\top}=-D$, if and only if the stencil satisfies
$w_{0}=0,\qquad w_{-\ell}=-w_{+\ell},\quad\ell=1,\dots,R.$ (10)
Proof.
By Lemma 1 the entries of $D$ are $D_{ij}=w_{i-j}$, so
$(D^{\top})_{ij}=D_{ji}=w_{j-i}.$
The condition $D^{\top}=-D$ is therefore equivalent to
[…]
or, writing $\ell=i-j$,
w=w+for all $_{-\ell}=-w_{+\ell}\quad\text${for all $\ell$.
In particular, taking $\ell=0$ yields $w_{0}=-w_{0}$, hence $w_{0}=0$, and for $\ell=1,\dots,R$ we obtain $w_{-\ell}=-w_{+\ell}$. Conversely, if (10) holds, then the above computation shows $D^{\top}=-D$, i.e., $D$ is skew-adjoint. ∎
We encode the conditions (10) as a set of linear constraints on the stencil vector $w$. Let $C\in\mathbb{R}^{(R+1)\times(2R+1)}$ be defined by
$(Cw)_{0}=w_{0},\qquad(Cw)_{\ell}=w_{-\ell}+w_{+\ell},\quad\ell=1,\dots,R,$
and let $d\in\mathbb{R}^{R+1}$ be the zero vector. Then the constraint $Cw=d$ is simply
$w_{0}=0,\qquad w_{-\ell}+w_{+\ell}=0,\quad\ell=1,\dots,R,$ (11)
which is equivalent to skew-symmetry.
Corollary 1.
Let $C$ and $d$ be defined in (11). Then $Cw=d$ if and only if the stencil $w$ satisfies (10), i.e., if and only if the corresponding convolution operator $D$ is skew-adjoint.
Proof.
Immediate from (11) and Theorem 2. ∎
Corollary 1 shows that the linear constraint $Cw=d$ used in our optimization problem is exactly the condition required for semi-discrete energy conservation via Theorem 1.
3 Wave speed and CFL for the learned operator
The discrete curl operator $D$ constructed in Section 2.1 is a periodic convolution with stencil $w=(w_{-R},\dots,w_{0},\dots,w_{R})$ satisfying the skew-symmetry constraints of Theorem 2. For such an operator it is convenient to work with its Fourier symbol
$\mu(\theta):=\sum_{k=-R}^{R}w_{k}\,\mathrm{e}^{\mathrm{i}k\theta},\qquad\theta\in[-\pi,\pi],$ (12)
which characterizes the response of $D$ to discrete Fourier modes $\mathrm{e}^{\mathrm{i}j\theta}$.
For the semi-discrete Maxwell system
$\frac{d}{dt}\begin{bmatrix}\mathbf{H}\\ \mathbf{E}\end{bmatrix}=A\begin{bmatrix}\mathbf{H}\\ \mathbf{E}\end{bmatrix},\qquad A=\begin{bmatrix}0&D\\ D^{T}&0\end{bmatrix}$,
the eigenvalues of $A$ are purely imaginary and can be written as $\lambda(\theta)=\pm\mathrm{i}\,|\mu(\theta)|$ for $\theta\in[-\pi,\pi]$. In particular, the quantity $|\mu(\theta)|$ plays the role of a numerical wave speed for the mode $\theta$, and the maximal value
$c_{\max}:=\max_{\theta\in[-\pi,\pi]}|\mu(\theta)|$ (13)
is the largest speed supported by the learned operator. For the standard centered difference stencil of width three one recovers $c_{\max}\sim h^{-1}$, corresponding to the usual CFL restriction for explicit wave propagation schemes.
If one were to apply the classical leapfrog scheme to the semi-discrete system, stability of the scalar test equation $\dot{u}=\lambda u$ with $\lambda=\mathrm{i}\omega$ requires $\Delta t\,|\omega|\leq 2$, see, e.g., [richtmyer1967, morton1994, leveque2007]. Substituting $\omega=\pm|\mu(\theta)|$ and maximizing over $\theta$ yields the CFL condition
$\Delta t\;\leq\;\frac{2}{c_{\max}},$ (14)
which expresses the admissible time step solely in terms of the learned symbol (12). In this way the learned stencil can be interpreted within the familiar CFL framework for explicit FDTD schemes, although in our numerical experiments we will use an implicit time integrator.
3.1 Crank–Nicolson time integration
For the actual time-dependent simulations in Section 6 we use the Crank–Nicolson (CN) method applied to the semi-discrete system
$\frac{d\mathbf{U}}{dt}=A\mathbf{U},\qquad\mathbf{U}=\begin{bmatrix}\mathbf{H}\\ \mathbf{E}\end{bmatrix},\qquad A^{T}=-A.$
The CN update reads
$\frac{\mathbf{U}^{n+1}-\mathbf{U}^{n}}{\Delta t}=A\,\frac{\mathbf{U}^{n+1}+\mathbf{U}^{n}}{2},$ (15)
which coincides with the implicit midpoint rule for linear systems. Since $A$ is skew-adjoint, the discrete energy
$\mathcal{E}^{n}:=\tfrac{1}{2}\|\mathbf{U}^{n}\|_{2}^{2}=\tfrac{1}{2}\bigl(\|\mathbf{H}^{n}\|_{2}^{2}+\|\mathbf{E}^{n}\|_{2}^{2}\bigr)$
is preserved exactly by (15) for any time step $\Delta t>0$; see, e.g., [hairer, hairerwanner2010] for background on such structure-preserving schemes.
In particular, CN does not require the CFL restriction (14) for stability or energy conservation in this linear, skew-adjoint setting. In the numerical experiments we nevertheless choose a time step $\Delta t$ proportional to the grid spacing $h$ and small enough that (14) would be satisfied for the corresponding centered-difference operator. This ensures that temporal discretization errors remain small relative to the spatial error while allowing us to interpret the learned stencil in terms of an effective wave speed $c_{\max}$.
4 Data-driven stencil learning
We now formulate the quadratic optimization problem used to learn stencils from data generated by a high-fidelity discretization.
4.1 Training data from spectral 1D Maxwell
We generate training data by sampling random Maxwell states and computing their time derivatives exactly (up to spectral discretization error). We work on the periodic domain $[0,L]$ with $N$ grid points and spacing $\Delta x=L/N$. For each training sample we draw a random truncated Fourier series
$E(x)$ $=\sum_{m=1}^{m_{\max}}a_{m}\sin\left(\frac{2\pi mx}{L}+\phi_{m}\right),$
$H(x)$ $=\sum_{m=1}^{m_{\max}}b_{m}\sin\left(\frac{2\pi mx}{L}+\psi_{m}\right),$
with amplitudes $a_{m},b_{m}$ and phases $\phi_{m},\psi_{m}$ drawn from simple Gaussian and uniform distributions. We then compute the spatial derivatives $\partial_{x}E$ and $\partial_{x}H$ spectrally using FFTs:
[…]
where $k$ is the discrete wavenumber. Finally, we obtain the time derivatives from the PDE (1):
$\partial_{t}E=\partial_{x}H,\qquad\partial_{t}H=\partial_{x}E.$
We denote by $U\in\mathbb{R}^{n_{\mathrm{sims}}\times 2\times N}$ the array of training states and by $dU/dt$ the corresponding array of time derivatives, with channel 0 equal to $E$ and channel $1$ equal to $H$.
4.2 Least-squares formulation
Given training pairs $(\mathbf{E},\mathbf{H})$ and $(\partial_{t}\mathbf{E},\partial_{t}\mathbf{H})$, we wish to learn a single convolution stencil $w$ such that
$D\mathbf{H}\approx\partial_{x}\mathbf{H}=\partial_{t}\mathbf{E},\qquad D\mathbf{E}\approx\partial_{x}\mathbf{E}=\partial_{t}\mathbf{H},$
in a least-squares sense. For each grid index $i$ away from the boundaries we extract local patches
$\mathbf{h}_{i}=(H_{i-R},\dots,H_{i+R})^{\top},\qquad\mathbf{e}_{i}=(E_{i-R},\dots,E_{i+R})^{\top},$
and define scalar targets
$y^{(H)}_{i}=\partial_{t}E_{i},\qquad y^{(E)}_{i}=\partial_{t}H_{i}.$
The desired relationships are
$w^{\top}\mathbf{h}_{i}\approx y^{(H)}_{i},\qquad w^{\top}\mathbf{e}_{i}\approx y^{(E)}_{i}$.
Stacking all patches and targets over space and all simulations, we obtain a linear system
$Aw\approx b,$ (16)
where each row of $A$ is either $\mathbf{h}_{i}^{\top}$ or $\mathbf{e}_{i}^{\top}$, and $b$ collects the corresponding targets. This leads to the unconstrained least-squares problem
$\min_{w\in\mathbb{R}^{2R+1}}\frac{1}{2}\|Aw-b\|_{2}^{2}+\frac{\lambda}{2}\|w\|_{2}^{2},$ (17)
where $\lambda\geq 0$ is a small Tikhonov regularization parameter. The function build_Ab_from_data in the code constructs $A$ and $b$.
4.3 Energy constraint and box bounds
To enforce skew-adjointness and hence discrete energy conservation, we impose the linear constraint $Cw=d$ with $C$ and $d$ defined as in Section 2.1. By Corollary 1, any stencil satisfying $Cw=d$ defines a skew-adjoint convolution operator $D$ and therefore an energy-conserving semi-discrete Maxwell system.
We also impose simple box bounds $|w_{\ell}|\leq M$, i.e.,
$M\leq w_{\ell}\leq M,\qquad\ell=-R,\dots,R,$ (18)
where $M>0$ is chosen large enough to contain the true derivative stencil but small enough to control numerical conditioning.
Our final optimization problem is the convex quadratic program
$\min_{w\in\mathbb{R}^{2R+1}}$ $\frac{1}{2}\|Aw-b\|_{2}^{2}+\frac{\lambda}{2}\|w\|_{2}^{2},$ (19)
s.t. $Cw=d,$
$-M\leq w\leq M.$
This problem is strictly convex for $\lambda>0$ and admits a unique solution. For our first-order methods we use a slightly simplified variant of (19) in which only the linear energy constraint is enforced. Specifically, projected gradient (PG) and Nesterov-accelerated gradient (NAG) are applied to the equality-constrained problem
$\min_{w\in\mathbb{R}^{2R+1}}$ $\frac{1}{2}\|Aw-b\|_{2}^{2}+\frac{\lambda}{2}\|w\|_{2}^{2},$ (20)
s.t. $Cw=d,$
while the full equality-plus-box-constrained problem (19) is handled by ADMM and the interior-point solver in CVXPY. In our experiments the solutions of (20) produced by PG and NAG remain well inside the box $M,M]^{2R+1}$, so the difference between (20) and (19) is negligible in practice.
5 Convex solvers
We briefly summarize the algorithms used to solve (19). All methods operate on the same design matrix $A$, constraint matrix $C$, and bounds.
5.1 Projected gradient methods
Our first two solvers are projected gradient (PG) and Nesterov-accelerated gradient (NAG), applied to the equality-constrained problem (20). In particular, for PG/NAG we enforce only the linear energy constraint $Cw=d$ and do not explicitly impose the box bounds $M\leq w\leq M$. In our experiments the PG/NAG iterates remain strictly inside the box, so the resulting solutions agree (to numerical tolerance) with those of the fully constrained problem (19).
Define the feasible affine subspace
$\mathcal{C}_{\mathrm{eq}}=\{w\in\mathbb{R}^{2R+1}:Cw=d\}.$
The orthogonal projection onto $\mathcal{C}_{\mathrm{eq}}$ with respect to the Euclidean inner product is given explicitly by
$P_{\mathcal{C}_{\mathrm{eq}}}(z):=z-C^{\top}(CC^{\top})^{-1}(Cz-d),\qquad z\in\mathbb{R}^{2R+1},$ (21)
assuming that $C$ has full row rank.
Let
$f(w):=\tfrac{1}{2}\|Aw-b\|_{2}^{2}+\tfrac{\lambda}{2}\|w\|_{2}^{2},\qquad\nabla f(w)=A^{\top}(Aw-b)+\lambda w.$
We choose a stepsize $\alpha=1/L$, where one may take $L=\|A\|_{2}^{2}+\lambda=\|A^{\top}A\|_{2}+\lambda$.
For PG, the update is
$w^{k+1}=P_{\mathcal{C}_{\mathrm{eq}}}\!\left(w^{k}-\alpha\nabla f(w^{k})\right),$
with $P_{\mathcal{C}_{\mathrm{eq}}}$ given by (21).
For NAG we use the extrapolated point
$y^{k}=w^{k}+\beta_{k}(w^{k}-w^{k-1}),$
and then update
$w^{k+1}=P_{\mathcal{C}_{\mathrm{eq}}}\!\left(y^{k}-\alpha\nabla f(y^{k})\right),$
with the standard parameters
$t_{k+1}=\tfrac{1}{2}\!\left(1+\sqrt{1+4t_{k}^{2}}\right),\qquad\beta_{k}=\frac{t_{k}-1}{t_{k+1}},\qquad t_{0}=1.$
We track the objective $f(w^{k})$, wall-clock time, and the equality residual $\|Cw^{k}-d\|_{2}$ at each iteration.
Remark 1.
Projecting sequentially onto $\{Cw=d\}$ and the box $M,M]^{2R+1}$ is not, in general, the exact Euclidean projection onto their intersection. If the box constraint becomes active and one wishes to enforce both constraints within a projected-gradient framework, one should compute $P_{\{Cw=d\}\cap[-M,M]^{2R+1}}$ using an intersection projection method such as Dykstra’s algorithm, or solve the projection step as a small quadratic program.
5.2 ADMM
To better handle the combination of equality and box constraints, we also implement an alternating direction method of multipliers (ADMM) scheme. We introduce a copy $z$ of $w$ and enforce the box constraint on $z$ only:
$w\in\mathbb{R}^{2R+1},\qquad z\in[-M,M]^{2R+1},\qquad w=z,\qquad Cw=d.$
The augmented Lagrangian splits into a quadratic $w$-problem with equality constraints and a simple proximal $z$-update given by box projection. The $w$-update reduces to solving a KKT system of the form
$\begin{bmatrix}A^{\top}A+(\lambda+\rho)I&C^{\top}\\ C&0\end{bmatrix}\begin{bmatrix}w\\ \eta\end{bmatrix}=\begin{bmatrix}A^{\top}b+\rho(z-u)\\ d\end{bmatrix}$,
where $\rho>0$ is the ADMM penalty parameter and $\eta$ is a vector of Lagrange multipliers for the equality constraints. The $z$-update is just componentwise clipping, $z^{k+1}=\Pi_{[-M,M]^{2R+1}}(w^{k+1}+u^{k})$, and the dual update is $u^{k+1}=u^{k}+(w^{k+1}-z^{k+1})$. We again monitor the objective, equality residual, and primal difference $\|w^{k+1}-w^{k}\|$.
5.3 Interior-point reference (CVXPY)
Finally, we use the CVXPY modeling framework with a cone solver (SCS in our experiments) to obtain a reference solution to (19). This interior-point-based solution serves as a baseline for the quality of the first-order methods. The corresponding routine solve_cvx sets up the quadratic objective and linear constraints, and returns the optimal stencil $w_{\mathrm{cvx}}$ when CVXPY is available.
6 Numerical experiments
We now present numerical experiments in the 1D Maxwell setting to evaluate the learned stencils and the performance of the different solvers. All experiments use the Crank–Nicolson (CN) time integrator with the learned spatial operator.
6.1 Experimental setup
Unless otherwise stated, we use the following default parameters:
  • Spatial domain $[0,L]$ with $L=1$ and $N=64$ cells, $\Delta x=L/N$.
  • Training data: $n_{\mathrm{sims}}=200$ random spectral samples with $m_{\max}=5$ Fourier modes.
  • Regularization parameter $\lambda=10^{-6}$ and box parameter $M=100$.
  • Stencil radii $R\in\{1,2,3,4\}$, corresponding to stencil length $2R+1$.
  • Time integration: Crank–Nicolson (CN) applied to the semi-discrete system with time step $\Delta t=0.5\Delta x$ and $N_{t}=300$ time steps. For skew-adjoint $D$, the CN update is norm-preserving, so no CFL restriction is required for stability, and the discrete energy is conserved up to roundoff.
The initial condition for time integration is chosen as a single Fourier mode,
$E(x,0)=\sin(2\pi x/L),\qquad H(x,0)=\cos(2\pi x/L),$
and we compute a reference solution using the exact central-difference stencil
$w_{\mathrm{exact}}=\frac{1}{2\Delta x}(-1,\,0,\,1).$
We then integrate the Maxwell system with each learned stencil using CN and measure the relative $L^{2}$ error in the electric field at final time:
$\mathrm{err}(w)=\frac{\|\mathbf{E}^{\mathrm{learned}}(T)-\mathbf{E}^{\mathrm{FD}}(T)\|_{2}}{\|\mathbf{E}^{\mathrm{FD}}(T)\|_{2}}.$ (22)
We also compute the energy-constraint residual
$r_{\mathrm{eq}}(w)=\|Cw-d\|_{2},$
which is zero for an exactly skew-symmetric stencil.
As an additional diagnostic, we track the discrete electromagnetic energy
$\mathcal{E}^{n}:=\frac{1}{2}\Delta x\sum_{i}\bigl[(E_{i}^{n})^{2}+(H_{i}^{n})^{2}\bigr]$
over time. For the exact stencil and CN time-stepping, the code confirms that $\mathcal{E}^{n}$ is constant to within floating-point roundoff, consistent with the analysis in Section 2.
6.2 Learned stencils for $R=1$
For $R=1$ the true centered-difference stencil (in normalized form) is $w_{\mathrm{exact}}=(-1,0,1)/(2\Delta x)$. Table 1 shows the learned stencils obtained by each solver compared with $w_{\mathrm{exact}}$. We observe that all methods recover a stencil very close to the exact one, with small differences due to the finite training set and regularization.
Method $w_{-1}$ $w_{0}$ $w_{+1}$
Exact CD 32.000000-32.000000 $0.000000$ $32.000000$
PG 32.906110-32.906110 $0.000000$ $32.906110$
NAG 32.906098-32.906098 $0.000000$ $32.906098$
ADMM 32.906114-32.906114 $2.64\!\times\!10^{-16}$ $32.906114$
CVX 32.906115-32.906115 $1.57\!\times\!10^{-14}$ $32.906115$
Table 1: Learned stencils for radius $R=1$. The exact centered difference (CD) stencil is compared against those learned by PG, NAG, ADMM, and CVX (values shown here are representative; see the accompanying code for the precise output on a given run).
Table 1 shows that all solvers recover a stencil very close to the exact centered difference. In Figure 1 we compare the final-time electric field $E(x,T)$ for each learned stencil against the reference FD solution; the curves lie almost on top of each other.
Figure 1: Final-time electric field $E(x,T)$ for the exact central-difference stencil and the learned stencils for $R=1$ under Crank–Nicolson time-stepping. The curves are visually indistinguishable at the scale shown.
Figure 2: Discrete energy error $E^{n}-E^{0}$ for the Crank–Nicolson time-stepping applied to the semi-discrete Maxwell system with a skew-symmetric discrete derivative $D$ (FD stencil, $R=1$). The error remains at the level of $10^{-14}$ over $N_{t}=300$ steps, consistent with conservation up to machine precision.
Figure 2 shows the discrete energy error $E^{n}-E^{0}$ over $N_{t}=300$ Crank–Nicolson steps. The drift remains at the level of $10^{-14}$, confirming that the fully discrete scheme preserves the semi-discrete energy up to roundoff.
6.3 Error and energy residual versus stencil radius
To explore the effect of stencil size, we repeat the learning and time-integration procedure for radii $R=1,2,3,4$. Figure 3 shows the final-time relative $L^{2}$ error $\mathrm{err}(w)$ for each method as a function of $R$.
Figure 3 shows that for $R=1$ all four solvers achieve comparable accuracy, with errors on the order of $5\times 10^{-3}$. As the radius increases, the behavior diverges: the projected-gradient and Nesterov methods exhibit substantially larger errors, with PG reaching $\mathcal{O}(1)$ error and NAG reaching $\mathcal{O}(10^{-2})$ for $R=3$. In contrast, the ADMM and CVX solutions maintain errors in a narrow band between approximately $3\times 10^{-3}$ and $6\times 10^{-3}$ for all radii tested. This indicates that the second-order KKT solve in ADMM and the interior-point method in CVX are more effective at exploiting the additional degrees of freedom introduced by larger stencils, whereas the first-order projected methods struggle to fully optimize the objective within the available iteration budget.
Figure 3: Final-time relative $L^{2}$ error in the electric field versus stencil radius $R$ for each solver, using Crank–Nicolson time integration.
Figure 4 reports the energy-constraint residual $r_{\mathrm{eq}}(w)=\lVert Cw-d\rVert_{2}$ as a function of the stencil radius $R$. The projected-gradient and Nesterov methods enforce the linear constraints to essentially machine precision for all radii, with residuals near $10^{-16}$, since each iteration includes an explicit orthogonal projection onto the affine subspace $\{w:Cw=d\}$. ADMM also attains very small residuals, in the range $10^{-15}$$10^{-13}$. The CVX solutions exhibit somewhat larger residuals, around $10^{-9}$$10^{-7}$, but these are still tiny on the scale of the stencil coefficients (which are $\mathcal{O}(10^{1})$). Thus all methods produce stencils that are effectively skew-symmetric, but ADMM and CVX combine this with significantly smaller PDE errors in Figure 3. In particular, these results show that driving $r_{\mathrm{eq}}(w)$ all the way to machine precision (as PG and NAG do) is not sufficient for accuracy if the least-squares objective is not simultaneously minimized; ADMM and CVX achieve a more favorable balance between data fit and constraint enforcement.
Figure 4: Energy-constraint residual $\|Cw-d\|_{2}$ versus stencil radius $R$ for each solver.
6.4 Solver convergence and runtime
Finally, we compare the convergence histories of the different solvers. Figure 5 plots the objective value versus iteration for each first-order method for a fixed radius. Figure 6 shows the same objective versus wall-clock time. These plots illustrate the tradeoffs among PG, NAG, and ADMM in terms of iterations, per-iteration cost, and constraint enforcement.
Figure 5: Objective value versus iteration for PG, NAG, and ADMM for a fixed radius $R$.
Figure 6: Objective value versus runtime for PG, NAG, and ADMM for a fixed radius $R$.
Figure 5 shows that all three methods reduce the objective by several orders of magnitude, but their convergence patterns differ. PG exhibits the expected monotone, roughly geometric decrease. NAG achieves a much faster reduction in the first $\sim 20$ iterations, but with the characteristic oscillations of accelerated gradient methods before settling near the optimum. ADMM appears as an almost flat curve at the bottom of the plot because it solves a small KKT system at each iteration; in this setting, one or two ADMM iterations are enough to reach an objective value comparable to the final PG/NAG iterates.
Figure 6 plots the same objective values against wall-clock time. Because PG and NAG have very cheap iterations, their curves almost coincide and reach a low objective well below $10^{-3}$ in about $10^{-2}$ seconds for our problem size. ADMM has a higher per-iteration cost due to the linear solve, but it reaches essentially its final objective in a single step, so the total runtime is competitive with the first-order methods in this regime. For larger-scale problems, we expect the trade-off between cheap but many iterations (PG/NAG) and expensive but very few iterations (ADMM) to be even more pronounced.
7 Conclusion
We proposed an energy-stable, data-driven discretization framework for the one-dimensional Maxwell system. By learning convolution stencils from spectral training data under a linear skew-symmetry constraint, we obtain discrete derivative operators that are guaranteed to be skew-adjoint and therefore exactly conserve semi-discrete electromagnetic energy. We demonstrated through numerical experiments that the resulting schemes match the accuracy of standard central differences while preserving structural properties. Using Crank–Nicolson time integration, we further confirmed that the discrete energy remains constant to roundoff for both hand-crafted and learned stencils satisfying the skew-adjoint constraint.
Although the present work focuses on a simple 1D model, the ideas are readily extensible to higher dimensions and more complex material models. Our results suggest that combining data-driven discretization with structure-preserving constraints is a promising direction for building stable, physically faithful reduced models in computational electromagnetics and beyond.
8 Discussion and extensions
The experiments above illustrate that simple linear constraints can encode important structural information in data-driven discretizations of PDEs. In this setting, the skew-symmetry constraint $Cw=d$ ensures semi-discrete energy conservation and leads to stencils that behave like standard centered differences but are obtained purely from data.
Future work will learn staggered curl operators that preserve a discrete Poynting theorem in 2D/3D, integrate material models (Lorentz/Duffing) with structure-preserving polarization updates, and incorporate boundary treatments (ABC/CPML). On the optimization side, extending the comparison to higher-order stencils, non-quadratic objectives, and learned nonlinear fluxes would help clarify the regimes in which ADMM or interior-point methods offer the best tradeoff relative to accelerated first-order algorithms. Ultimately, combining learning with physics-motivated constraints promises accurate, stable reduced models for large-scale CEM while keeping the mathematical guarantees that make FDTD dependable [taf-love, hairer].

### v2

@@
 1 Introduction
-Finite-difference time-domain (FDTD) methods are a standard tool for solving Maxwell’s equations in computational electromagnetics [taf-love].
+Finite-difference time-domain (FDTD) methods are a standard tool for solving Maxwell’s equations in computational electromagnetics [15].
 Classical schemes, such as the Yee method, are derived by hand from the continuous equations, with spatial derivatives approximated using fixed central-difference stencils and stability enforced via analytical CFL conditions.
 Recently, there has been growing interest in data-driven discretizations, where discrete operators are learned from numerical or experimental data, e.g., via regression or neural networks.
-Examples include data-driven discretization of Burgers and Navier–Stokes equations and learned coarse-grid operators for PDEs [bar-sinai].
+Examples include data-driven discretization of Burgers and Navier–Stokes equations and learned coarse-grid operators for PDEs [1].
 Most existing data-driven schemes aim to minimize approximation error but do not explicitly enforce structural properties such as stability, conservation, or energy balance.
@@
 There is a growing literature on learning discretizations and coarse-grid operators for PDEs directly from data.
-A prominent example is the work of [bar-sinai], who learn finite-difference schemes for Burgers and Navier–Stokes equations by fitting local stencils to high-fidelity solutions while optimizing accuracy on a training set.
-Related approaches use regression or neural networks to construct reduced models or learned fluxes for time-dependent PDEs; see, e.g., [parish2016paradigm, ling2016reynolds, rudy2017data, raissi2018hidden, kochkov2021machine].
+A prominent example is the work of [1], who learn finite-difference schemes for Burgers and Navier–Stokes equations by fitting local stencils to high-fidelity solutions while optimizing accuracy on a training set.
+Related approaches use regression or neural networks to construct reduced models or learned fluxes for time-dependent PDEs; see, e.g., [11, 9, 14, 12, 7].
 In most of these methods, the learned operators are chosen solely to minimize an empirical error measure and are not required to satisfy structural constraints such as conservation, stability, or symmetries of the underlying PDE.
@@
 Energy-preserving FDTD and skew-adjoint schemes.
-Classical finite-difference time-domain (FDTD) methods for Maxwell’s equations, such as the Yee scheme and its variants, are carefully constructed so that discrete versions of divergence constraints and energy identities hold at the semi-discrete or fully discrete level [taf-love, richtmyer1967, morton1994, leveque2007].
-In particular, staggered spatial discretizations and symplectic or A-stable time integrators such as the Crank–Nicolson scheme are often used to obtain skew-adjoint or nearly skew-adjoint update operators, leading to strong stability and conservation properties [crank1947, giles2010, hairerwanner2010].
-There is also a large body of work on structure-preserving and geometric integrators for Hamiltonian ODE/PDE systems, which emphasizes invariants such as energy and symplectic form [hairer].
+Classical finite-difference time-domain (FDTD) methods for Maxwell’s equations, such as the Yee scheme and its variants, are carefully constructed so that discrete versions of divergence constraints and energy identities hold at the semi-discrete or fully discrete level [15, 13, 10, 8].
+In particular, staggered spatial discretizations and symplectic or A-stable time integrators such as the Crank–Nicolson scheme are often used to obtain skew-adjoint or nearly skew-adjoint update operators, leading to strong stability and conservation properties [3, 4, 6].
+There is also a large body of work on structure-preserving and geometric integrators for Hamiltonian ODE/PDE systems, which emphasizes invariants such as energy and symplectic form [5].
 Our work is closest in spirit to these structure-preserving schemes, but differs in that the discrete spatial operator is learned from data rather than hand-designed: we encode skew-adjointness as a set of linear constraints on the stencil coefficients and solve a convex quadratic program to obtain an energy-conserving discretization.
@@
 $(D\mathbf{u})_{i}=\sum_{\ell=-R}^{R}w_{\ell}\,u_{i+\ell}=\sum_{j=0}^{N-1}w_{i-j}\,u_{j},$
-where $j=i+\ell$ (modulo $N$). Hence $D_{ij}=w_{i-j}$. ∎
+where $j=i+\ell$ (modulo NN). Hence $D_{ij}=w_{i-j}$.[2]
 Theorem 2 (Characterization of skew-symmetric stencils).
-Let $D$ be the convolution operator (2) with periodic boundary conditions and stencil $w\in\mathbb{R}^{2R+1}$.
+[2]Let $D$ be the convolution operator (2) with periodic boundary conditions and stencil $w\in\mathbb{R}^{2R+1}$.
 Then $D$ is skew-adjoint with respect to (4), i.e., $D^{\top}=-D$, if and only if the stencil satisfies
@@
 For the standard centered difference stencil of width three one recovers $c_{\max}\sim h^{-1}$, corresponding to the usual CFL restriction for explicit wave propagation schemes.
-If one were to apply the classical leapfrog scheme to the semi-discrete system, stability of the scalar test equation $\dot{u}=\lambda u$ with $\lambda=\mathrm{i}\omega$ requires $\Delta t\,|\omega|\leq 2$, see, e.g., [richtmyer1967, morton1994, leveque2007].
+If one were to apply the classical leapfrog scheme to the semi-discrete system, stability of the scalar test equation $\dot{u}=\lambda u$ with $\lambda=\mathrm{i}\omega$ requires $\Delta t\,|\omega|\leq 2$, see, e.g., [13, 10, 8].
 Substituting $\omega=\pm|\mu(\theta)|$ and maximizing over $\theta$ yields the CFL condition
@@
 $\mathcal{E}^{n}:=\tfrac{1}{2}\|\mathbf{U}^{n}\|_{2}^{2}=\tfrac{1}{2}\bigl(\|\mathbf{H}^{n}\|_{2}^{2}+\|\mathbf{E}^{n}\|_{2}^{2}\bigr)$
-is preserved exactly by (15) for any time step $\Delta t>0$; see, e.g., [hairer, hairerwanner2010] for background on such structure-preserving schemes.
+is preserved exactly by (15) for any time step $\Delta t>0$; see, e.g., [5, 6] for background on such structure-preserving schemes.
 In particular, CN does not require the CFL restriction (14) for stability or energy conservation in this linear, skew-adjoint setting.
@@
 $\partial_{t}E=\partial_{x}H,\qquad\partial_{t}H=\partial_{x}E.$
-We denote by $U\in\mathbb{R}^{n_{\mathrm{sims}}\times 2\times N}$ the array of training states and by $dU/dt$ the corresponding array of time derivatives, with channel 0 equal to $E$ and channel $1$ equal to $H$.
+We denote by
+$U\in\mathbb{R}^{n_{\mathrm{sims}}\times 2\times N}$
+the array of training states, where $U[s,0,i]=E^{(s)}(x_{i})$ and $U[s,1,i]=H^{(s)}(x_{i})$ denote the electric and magnetic fields of the $s$-th training sample evaluated at the grid point $x_{i}$.
+Similarly,
+$\frac{dU}{dt}\in\mathbb{R}^{n_{\mathrm{sims}}\times 2\times N}$
+stores the corresponding time derivatives, with $(dU/dt)[s,0,i]=\partial_{t}E^{(s)}(x_{i})$ and $(dU/dt)[s,1,i]=\partial_{t}H^{(s)}(x_{i})$.
+Here $n_{\mathrm{sims}}$ denotes the number of independent training samples (or snapshots), each consisting of a randomly generated Maxwell state $(E,H)$ and its corresponding time derivative.
+The spatial domain $[0,L]$ is discretized using $N$ uniformly spaced grid points $x_{i}=i\Delta x$, $i=0,\dots,N-1$, where $\Delta x=L/N$.
 4.2 Least-squares formulation
@@
 $\min_{w\in\mathbb{R}^{2R+1}}\frac{1}{2}\|Aw-b\|_{2}^{2}+\frac{\lambda}{2}\|w\|_{2}^{2},$ (17)
-where $\lambda\geq 0$ is a small Tikhonov regularization parameter. The function build_Ab_from_data in the code constructs $A$ and $b$.
+where $\lambda\geq 0$ is a small Tikhonov regularization parameter.
 4.3 Energy constraint and box bounds
@@
   • Time integration: Crank–Nicolson (CN) applied to the semi-discrete system with time step $\Delta t=0.5\Delta x$ and $N_{t}=300$ time steps.
-For skew-adjoint $D$, the CN update is norm-preserving, so no CFL restriction is required for stability, and the discrete energy is conserved up to roundoff.
 The initial condition for time integration is chosen as a single Fourier mode,
@@
 and we compute a reference solution using the exact central-difference stencil
-$w_{\mathrm{exact}}=\frac{1}{2\Delta x}(-1,\,0,\,1).$
+$w_{\mathrm{exact}}=\frac{1}{2\Delta x}(-1,\,0,\,1),$ (22)
 We then integrate the Maxwell system with each learned stencil using CN and measure the relative $L^{2}$ error in the electric field at final time:
@@
 6.2 Learned stencils for $R=1$
-For $R=1$ the true centered-difference stencil (in normalized form) is $w_{\mathrm{exact}}=(-1,0,1)/(2\Delta x)$.
+For $R=1$ the true centered-difference stencil is precisely $w_{\mathrm{exact}}$ from (22).
 Table 1 shows the learned stencils obtained by each solver compared with $w_{\mathrm{exact}}$.
@@
 Method $w_{-1}$ $w_{0}$ $w_{+1}$
-Exact CD 32.000000-32.000000 $0.000000$ $32.000000$
+Exact FD 32.000000-32.000000 $0.000000$ $32.000000$
 PG 32.906110-32.906110 $0.000000$ $32.906110$
@@
 Table 1: Learned stencils for radius $R=1$.
-The exact centered difference (CD) stencil is compared against those learned by PG, NAG, ADMM, and CVX (values shown here are representative; see the accompanying code for the precise output on a given run).
 Table 1 shows that all solvers recover a stencil very close to the exact centered difference.
-In Figure 1 we compare the final-time electric field $E(x,T)$ for each learned stencil against the reference FD solution; the curves lie almost on top of each other.
+In In Figures 1 and 2 we compare the space–time evolution and the final-time profile of the electric field $E(x,t)$ for each learned stencil against the reference FD solution.
+The space–time plots in Figure 1 show that the learned stencils reproduce essentially the same wave propagation as the exact centered difference, while the final-time profiles in Figure 2 are visually indistinguishable from the reference solution.
+Figure 1: Space–time evolution of the electric field $E(x,t)$ for the reference central-difference stencil and the learned stencils with radius $R=1$.
+Each panel shows $E(x,t)$ over the full simulation time.
 Figure 2: Final-time electric field $E(x,T)$ for the exact central-difference stencil and the learned stencils for $R=1$ under Crank–Nicolson time-stepping.
 The curves are visually indistinguishable at the scale shown.
-Figure 2: Discrete energy error $E^{n}-E^{0}$ for the Crank–Nicolson time-stepping applied to the semi-discrete Maxwell system with a skew-symmetric discrete derivative $D$ (FD stencil, $R=1$).
-The error remains at the level of $10^{-14}$ over $N_{t}=300$ steps, consistent with conservation up to machine precision.
-Figure 2 shows the discrete energy error $E^{n}-E^{0}$ over $N_{t}=300$ Crank–Nicolson steps.
-The drift remains at the level of $10^{-14}$, confirming that the fully discrete scheme preserves the semi-discrete energy up to roundoff.
 6.3 Error and energy residual versus stencil radius
@@
 Figure 3: Final-time relative $L^{2}$ error in the electric field versus stencil radius $R$ for each solver, using Crank–Nicolson time integration.
-Figure 4 reports the energy-constraint residual $r_{\mathrm{eq}}(w)=\lVert Cw-d\rVert_{2}$ as a function of the stencil radius $R$.
-The projected-gradient and Nesterov methods enforce the linear constraints to essentially machine precision for all radii, with residuals near $10^{-16}$, since each iteration includes an explicit orthogonal projection onto the affine subspace $\{w:Cw=d\}$.
-ADMM also attains very small residuals, in the range $10^{-15}$$10^{-13}$.
-The CVX solutions exhibit somewhat larger residuals, around $10^{-9}$$10^{-7}$, but these are still tiny on the scale of the stencil coefficients (which are $\mathcal{O}(10^{1})$).
-Thus all methods produce stencils that are effectively skew-symmetric, but ADMM and CVX combine this with significantly smaller PDE errors in Figure 3.
-In particular, these results show that driving $r_{\mathrm{eq}}(w)$ all the way to machine precision (as PG and NAG do) is not sufficient for accuracy if the least-squares objective is not simultaneously minimized; ADMM and CVX achieve a more favorable balance between data fit and constraint enforcement.
-Figure 4: Energy-constraint residual $\|Cw-d\|_{2}$ versus stencil radius $R$ for each solver.
+6.4 Discrete energy conservation for learned Maxwell stencils
+We next examine how closely the learned stencils preserve the semi-discrete electromagnetic energy under Crank–Nicolson time integration.
+Figure 4: Discrete energy error $\mathcal{E}^{n}-\mathcal{E}^{0}$ for Crank–Nicolson time stepping with the exact central-difference stencil and the structure-preserving learned stencils, shown on a linear (left) and logarithmic (right) scale.
+Figure 4 compares the discrete electromagnetic energy error $\mathcal{E}^{n}-\mathcal{E}^{0}$ for Crank–Nicolson time integration using the exact central-difference stencil and the structure-preserving learned stencils.
+On the linear scale, all curves are essentially flat and remain at the level of $10^{-12}$, indicating that the total energy is conserved up to machine precision for all methods.
+The logarithmic view reveals small differences: the PG, ADMM, and CVX stencils produce energy drift of the same order as the exact FD stencil, while NAG exhibits a slightly larger but still very small drift.
+These results confirm that enforcing the skew-symmetry constraints during training yields learned stencils that are effectively energy-preserving under Crank–Nicolson time stepping.
 6.5 Solver convergence and runtime
 Finally, we compare the convergence histories of the different solvers.
-Figure 5 plots the objective value versus iteration for each first-order method for a fixed radius.
-Figure 6 shows the same objective versus wall-clock time.
+Figure 5(a) plots the objective value versus iteration for each first-order method for a fixed radius.
+Figure 5(b) shows the same objective versus wall-clock time.
 These plots illustrate the tradeoffs among PG, NAG, and ADMM in terms of iterations, per-iteration cost, and constraint enforcement.
-Figure 5: Objective value versus iteration for PG, NAG, and ADMM for a fixed radius $R$.
-Figure 6: Objective value versus runtime for PG, NAG, and ADMM for a fixed radius $R$.
-Figure 5 shows that all three methods reduce the objective by several orders of magnitude, but their convergence patterns differ.
+(a) Objective vs iteration
+(b) Objective vs runtime
+Figure 5: Convergence behavior of PG, NAG, and ADMM for a fixed stencil radius $R$. Left: objective value versus iteration count. Right: objective value versus wall-clock runtime.
+Figure 5(a) shows that all three methods reduce the objective by several orders of magnitude, but their convergence patterns differ.
 PG exhibits the expected monotone, roughly geometric decrease.
@@
 ADMM appears as an almost flat curve at the bottom of the plot because it solves a small KKT system at each iteration; in this setting, one or two ADMM iterations are enough to reach an objective value comparable to the final PG/NAG iterates.
-Figure 6 plots the same objective values against wall-clock time.
+Figure 5(b) plots the same objective values against wall-clock time.
 Because PG and NAG have very cheap iterations, their curves almost coincide and reach a low objective well below $10^{-3}$ in about $10^{-2}$ seconds for our problem size.
@@
 7 Conclusion
-We proposed an energy-stable, data-driven discretization framework for the one-dimensional Maxwell system.
+We proposed an energy-conserving, data-driven discretization framework for the one-dimensional Maxwell system.
 By learning convolution stencils from spectral training data under a linear skew-symmetry constraint, we obtain discrete derivative operators that are guaranteed to be skew-adjoint and therefore exactly conserve semi-discrete electromagnetic energy.
-We demonstrated through numerical experiments that the resulting schemes match the accuracy of standard central differences while preserving structural properties.
-Using Crank–Nicolson time integration, we further confirmed that the discrete energy remains constant to roundoff for both hand-crafted and learned stencils satisfying the skew-adjoint constraint.
-Although the present work focuses on a simple 1D model, the ideas are readily extensible to higher dimensions and more complex material models.
-Our results suggest that combining data-driven discretization with structure-preserving constraints is a promising direction for building stable, physically faithful reduced models in computational electromagnetics and beyond.
+Numerical experiments show that the resulting schemes match the accuracy of standard central differences while preserving these structural properties.
+Using Crank–Nicolson time integration, we further confirmed that the discrete energy remains constant to roundoff for learned stencils satisfying the skew-adjoint constraint.
 8 Discussion and extensions
@@
 On the optimization side, extending the comparison to higher-order stencils, non-quadratic objectives, and learned nonlinear fluxes would help clarify the regimes in which ADMM or interior-point methods offer the best tradeoff relative to accelerated first-order algorithms.
-Ultimately, combining learning with physics-motivated constraints promises accurate, stable reduced models for large-scale CEM while keeping the mathematical guarantees that make FDTD dependable [taf-love, hairer].
+Ultimately, combining learning with physics-motivated constraints promises accurate, stable reduced models for large-scale CEM while keeping the mathematical guarantees that make FDTD dependable [15, 5].

### v3

@@
-Energy-Conserving Data-Driven Discretizations for Maxwell’s Equations
+Physics-Constrained Learning of Energy-Preserving Stencils for Maxwell’s Equations
 Abstract
 We study data-driven construction of spatial discretizations for the one-dimensional Maxwell system.
-Given high-fidelity training data generated by a spectral discretization, we learn a linear convolution stencil that approximates the spatial derivative operator appearing in Maxwell’s equations.
-The stencil is obtained by solving a convex quadratic optimization problem, subject to linear constraints that enforce skew-adjointness of the discrete derivative.
-These constraints guarantee a semi-discrete energy identity for the resulting Maxwell system.
-We prove that our constraints characterize the class of skew-symmetric convolution operators and express the associated numerical wave speed and CFL restriction for the classical leapfrog scheme in terms of the learned stencil’s Fourier symbol.
-We then compare several convex solvers for the resulting quadratic program—projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point reference implemented in CVXPY—and evaluate the learned schemes in time-dependent one-dimensional Maxwell simulations using a Crank–Nicolson (CN) time discretization.
-Our numerical experiments show that (i) energy-constrained learned stencils achieve accuracy comparable to standard central differences while exactly preserving the discrete electromagnetic energy under CN time-stepping, and (ii) ADMM and interior-point methods produce nearly identical operators, with ADMM offering a favorable tradeoff between accuracy, constraint satisfaction, and runtime.
+Using high-fidelity training data from a spectral discretization, we learn a linear convolution stencil that approximates the spatial derivative operator in Maxwell’s equations.
+We formulate a convex quadratic program for the stencil coefficients with linear constraints that enforce skew-adjointness of the discrete derivative; these constraints guarantee a semi-discrete electromagnetic energy identity and yield a CFL condition expressed directly in terms of the stencil’s Fourier symbol.
+We compare several convex solvers for the resulting quadratic program—projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point reference implemented in CVXPY—and evaluate the learned operators in time-dependent Maxwell simulations using a Crank–Nicolson (CN) discretization.
+Numerical experiments, including cases with nonstandard target operators and noisy training data, show that (i) energy-constrained learned stencils achieve accuracy comparable to standard central differences while exactly preserving the discrete electromagnetic energy under CN time-stepping, and (ii) ADMM and interior-point solvers produce nearly identical operators, with ADMM offering a favorable tradeoff between accuracy, constraint satisfaction, and runtime.
 1 Introduction
@@
 Classical schemes, such as the Yee method, are derived by hand from the continuous equations, with spatial derivatives approximated using fixed central-difference stencils and stability enforced via analytical CFL conditions.
-Recently, there has been growing interest in data-driven discretizations, where discrete operators are learned from numerical or experimental data, e.g., via regression or neural networks.
+Recently, there has been growing interest in data-driven discretizations, where discrete operators are learned from numerical or experimental data, for example via regression or neural networks.
 Examples include data-driven discretization of Burgers and Navier–Stokes equations and learned coarse-grid operators for PDEs [1].
@@
 For Maxwell’s equations this is a significant limitation, since standard FDTD schemes are carefully designed so that discrete versions of divergence constraints and energy identities hold at the semi-discrete level.
-In this work we address this gap in a simple but concrete setting: the one-dimensional TE Maxwell system with periodic boundary conditions. Our goal is to learn a discrete spatial derivative operator
+In this work we address this gap in a simple but concrete setting: the one-dimensional Maxwell system with periodic boundary conditions. Our goal is to learn a discrete spatial derivative operator
 $D:\mathbb{R}^{N}\to\mathbb{R}^{N}$
@@
 This structural constraint guarantees exact conservation of discrete electromagnetic energy at the semi-discrete level, regardless of the details of the training data.
-Novelty and contributions.
-Building on this context, our main contributions are as follows:
-  • We formulate a convex quadratic optimization problem for learning convolution stencils for the 1D Maxwell system from spectral training data.
-The objective is a least-squares misfit between the learned discrete derivative and high-fidelity derivatives, while the constraints encode skew-symmetry of the stencil coefficients together with simple box bounds.
-  • We show that these linear constraints are exactly equivalent to skew-adjointness of the corresponding convolution operator with respect to the discrete $L^{2}$ inner product, and hence to a semi-discrete electromagnetic energy identity.
-This yields, to our knowledge, the first learned Maxwell stencil with a guaranteed semi-discrete energy conservation law.
-  • We characterize the numerical wave speed and associated CFL restriction of the learned discrete curl operator purely in terms of its Fourier symbol $\mu(\theta)$.
-  • We implement projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point method (via CVXPY) to solve the constrained quadratic program, and we benchmark the resulting stencils in time-dependent 1D Maxwell simulations.
-The experiments compare PDE error, energy-constraint residuals, and energy evolution across stencil radii, and demonstrate that ADMM and interior-point methods produce nearly identical, highly accurate stencils, with ADMM offering a favorable tradeoff between accuracy, constraint satisfaction, and runtime.
+Classical FDTD schemes such as the Yee method [21] are hand-derived so that the discrete curl operator is skew-adjoint, yielding exact semi-discrete energy conservation.
+High-order extensions and operator-splitting schemes similarly maintain this structure [3, 6].
+Data-driven PDE schemes learn local stencils or coarse-grid operators directly from numerical data.
+Bar-Sinai et al.
+learn finite-difference discretizations for Burgers and Navier–Stokes, showing that learned operators can match or exceed classical schemes on specific training distributions [1].
+Many follow-up works learn PDE operators or closure models [17, 11, 20, 18, 9], but typically do not enforce stability, skew-adjointness, or conservation laws.
+Recent neural-operator surrogates for electromagnetics incorporate physics through architectural biases or penalty terms.
+Dynamic-convolution neural operators for FDTD-like EM propagation impose causal receptive fields and material dispersion structure [14].
+Domain-decomposition neural operators couple learned subdomain solvers through interface penalties or consistency conditions [15, 8].
+Other neural operator frameworks embed conservation via soft constraints or corrective layers [13, 12].
+These approaches improve stability but enforce physical structure only approximately, through optimization.
+Building on this context, our main contributions are as follows.
+First, we formulate a convex quadratic optimization problem for learning convolution stencils for the 1D Maxwell system from spectral training data, with an objective given by a least-squares misfit between the learned discrete derivative and high-fidelity derivatives, and linear constraints that encode skew-symmetry of the stencil coefficients together with simple box bounds.
+Second, we show that these linear constraints are exactly equivalent to skew-adjointness of the corresponding convolution operator with respect to the discrete $L^{2}$ inner product, and hence to a semi-discrete electromagnetic energy identity, yielding what is, to our knowledge, the first learned Maxwell stencil with a guaranteed semi-discrete energy conservation law.
+Third, we characterize the numerical wave speed and associated CFL restriction of the learned discrete curl operator purely in terms of its Fourier symbol $\mu(\theta)$.
+Finally, we implement projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point method (via CVXPY) to solve the constrained quadratic program, and we benchmark the resulting stencils in time-dependent 1D Maxwell simulations; the experiments compare PDE error, energy-constraint residuals, and energy evolution across stencil radii, and demonstrate that ADMM and interior-point methods produce nearly identical, highly accurate stencils, with ADMM offering a favorable tradeoff between accuracy, constraint satisfaction, and runtime.
 Although our experiments focus on a simple 1D periodic model, the framework and energy constraints extend naturally to higher-dimensional settings and more complex material models.
 Our results illustrate how even simple linear constraints in a data-driven discretization can encode important physics and restore structural properties such as energy conservation.
-1.1 Related work
-Data-driven discretizations.
-There is a growing literature on learning discretizations and coarse-grid operators for PDEs directly from data.
-A prominent example is the work of [1], who learn finite-difference schemes for Burgers and Navier–Stokes equations by fitting local stencils to high-fidelity solutions while optimizing accuracy on a training set.
-Related approaches use regression or neural networks to construct reduced models or learned fluxes for time-dependent PDEs; see, e.g., [11, 9, 14, 12, 7].
-In most of these methods, the learned operators are chosen solely to minimize an empirical error measure and are not required to satisfy structural constraints such as conservation, stability, or symmetries of the underlying PDE.
-In contrast, our focus is on learning a discrete Maxwell operator that provably satisfies a semi-discrete energy identity through an explicit skew-adjointness constraint on the convolution stencil.
-Energy-preserving FDTD and skew-adjoint schemes.
-Classical finite-difference time-domain (FDTD) methods for Maxwell’s equations, such as the Yee scheme and its variants, are carefully constructed so that discrete versions of divergence constraints and energy identities hold at the semi-discrete or fully discrete level [15, 13, 10, 8].
-In particular, staggered spatial discretizations and symplectic or A-stable time integrators such as the Crank–Nicolson scheme are often used to obtain skew-adjoint or nearly skew-adjoint update operators, leading to strong stability and conservation properties [3, 4, 6].
-There is also a large body of work on structure-preserving and geometric integrators for Hamiltonian ODE/PDE systems, which emphasizes invariants such as energy and symplectic form [5].
-Our work is closest in spirit to these structure-preserving schemes, but differs in that the discrete spatial operator is learned from data rather than hand-designed: we encode skew-adjointness as a set of linear constraints on the stencil coefficients and solve a convex quadratic program to obtain an energy-conserving discretization.
-This work in context.
-To the best of our knowledge, the present paper is the first to combine data-driven stencil learning for Maxwell’s equations with linear constraints that (i) exactly characterize skew-symmetric convolution operators and (ii) yield a rigorous semi-discrete energy identity and a CFL condition expressed directly in terms of the learned stencil’s Fourier symbol.
+Our method imposes hard linear constraints that make the learned convolution stencil exactly skew-adjoint, guaranteeing a discrete electromagnetic energy identity independent of training accuracy or solver choice.
+This places our work closest to Bar-Sinai et al.
+but with strict structure preservation rather than penalty-based regularization, and complementary to physics-informed neural operators, which typically enforce conservation approximately.
+To the best of our knowledge, this is the first work to combine data-driven stencil learning for Maxwell’s equations with linear constraints that (i) exactly characterize skew-symmetric convolution operators and (ii) yield a rigorous semi-discrete energy identity and a CFL condition expressed directly in terms of the learned stencil’s Fourier symbol.
 This provides a bridge between unconstrained data-driven discretizations and classical energy-preserving FDTD schemes, while remaining compatible with standard convex optimization toolchains (projected gradient, Nesterov acceleration, ADMM, and interior-point solvers).
+Beyond the homogeneous model, Sections 6.6 and 6.7 illustrate how the same constrained learning framework can recover nonstandard, effective operators and regularize noisy or mildly inconsistent training data.
 2 Semi-discrete Maxwell system and energy identity
@@
 which expresses the admissible time step solely in terms of the learned symbol (12).
-In this way the learned stencil can be interpreted within the familiar CFL framework for explicit FDTD schemes, although in our numerical experiments we will use an implicit time integrator.
+In this way the learned stencil can be interpreted within the familiar CFL framework for explicit FDTD schemes, although in our numerical experiments we will use an implicit time integrator.In this work we focus on CN, which is unconditionally stable for the skew-adjoint semi-discrete operator.
+A systematic comparison with leapfrog using the derived CFL bound is left for future work
 3.1 Crank–Nicolson time integration
@@
 $D\mathbf{H}\approx\partial_{x}\mathbf{H}=\partial_{t}\mathbf{E},\qquad D\mathbf{E}\approx\partial_{x}\mathbf{E}=\partial_{t}\mathbf{H},$
-in a least-squares sense. For each grid index $i$ away from the boundaries we extract local patches
+in a least-squares sense. Because we assume periodic boundary conditions, the stencil is applied identically at every grid point, with indices interpreted modulo $N$.
+For each grid index $i$ we extract local patches
 $\mathbf{h}_{i}=(H_{i-R},\dots,H_{i+R})^{\top},\qquad\mathbf{e}_{i}=(E_{i-R},\dots,E_{i+R})^{\top},$
@@
 In our experiments the solutions of (20) produced by PG and NAG remain well inside the box $M,M]^{2R+1}$, so the difference between (20) and (19) is negligible in practice.
+In the presence of noisy or mildly inconsistent training data, the least-squares system can become ill-conditioned and produce stencils with very large coefficients; the Tikhonov term and box bounds mitigate this behavior, as illustrated in Section 6.7.
 5 Convex solvers
@@
 5.2 ADMM
-To better handle the combination of equality and box constraints, we also implement an alternating direction method of multipliers (ADMM) scheme.
+To better handle the combination of equality and box constraints, we also implement an alternating direction method of multipliers (ADMM) [4], [5] scheme.
 We introduce a copy $z$ of $w$ and enforce the box constraint on $z$ only:
@@
 6 Numerical experiments
+All experiments were implemented in Python using NumPy for linear algebra and SciPy for FFTs.
+The ADMM and projected-gradient/Nesterov solvers are written explicitly as described in Section 5.
+The interior-point reference solution is computed using CVXPY (version X.Y) with the SCS solver.
+Runtimes are reported for a single-threaded run on a (Apple M4, 16 GB RAM) using Python 3.11.
+Although absolute times depend on hardware and implementation details, the relative trends between first-order methods (PG/NAG) and the second-order ADMM/CVX approach remained consistent across several runs.
 We now present numerical experiments in the 1D Maxwell setting to evaluate the learned stencils and the performance of the different solvers.
@@
 6.2 Learned stencils for $R=1$
-For $R=1$ the true centered-difference stencil is precisely $w_{\mathrm{exact}}$ from (22).
 Table 1 shows the learned stencils obtained by each solver compared with $w_{\mathrm{exact}}$.
@@
 Table 1: Learned stencils for radius $R=1$.
-Table 1 shows that all solvers recover a stencil very close to the exact centered difference.
-In In Figures 1 and 2 we compare the space–time evolution and the final-time profile of the electric field $E(x,t)$ for each learned stencil against the reference FD solution.
-The space–time plots in Figure 1 show that the learned stencils reproduce essentially the same wave propagation as the exact centered difference, while the final-time profiles in Figure 2 are visually indistinguishable from the reference solution.
-Figure 1: Space–time evolution of the electric field $E(x,t)$ for the reference central-difference stencil and the learned stencils with radius $R=1$.
-Each panel shows $E(x,t)$ over the full simulation time.
+In In Figures 1 and 2, we compare the space-time evolution and the final-time profile of the electric field $E(x,t)$ for each learned stencil against the reference FD solution for R =1.
+The final-time and space-time profiles in Figure 2 are visually indistinguishable from the reference solution.
+Figure 1: Space-Time Evolution
 Figure 2: Final-time electric field $E(x,T)$ for the exact central-difference stencil and the learned stencils for $R=1$ under Crank–Nicolson time-stepping.
 The curves are visually indistinguishable at the scale shown.
+This experiment serves as a sanity check for the learning framework: for the smallest radius $R=1$, the energy-constrained optimization recovers a stencil that is essentially indistinguishable from the classical centered difference, and the corresponding Crank–Nicolson time integration reproduces the reference solution to within plotting accuracy.
+In other words, when the true operator is known and low order, the learned discretizations do not degrade either accuracy or energy behavior, while providing the same skew-adjoint structure that underpins exact discrete energy conservation.
+This $R=1$ case therefore provides a baseline for the higher-radius stencils and noisy-training experiments considered below.
 6.3 Error and energy residual versus stencil radius
@@
 Figure 3 shows the final-time relative $L^{2}$ error $\mathrm{err}(w)$ for each method as a function of $R$.
+We initialize all optimizations at the zero stencil; even from this uninformative start, the energy-constrained learner recovers stencils close to classical centered differences.
 Figure 3 shows that for $R=1$ all four solvers achieve comparable accuracy, with errors on the order of $5\times 10^{-3}$.
@@
 In contrast, the ADMM and CVX solutions maintain errors in a narrow band between approximately $3\times 10^{-3}$ and $6\times 10^{-3}$ for all radii tested.
-This indicates that the second-order KKT solve in ADMM and the interior-point method in CVX are more effective at exploiting the additional degrees of freedom introduced by larger stencils, whereas the first-order projected methods struggle to fully optimize the objective within the available iteration budget.
+This indicates that the second-order KKT solve in ADMM and the interior-point method in CVX are more effective at exploiting the additional degrees of freedom introduced by larger stencils, whereas the first-order projected methods struggle to fully optimize the objective within the available iteration budget.We believe the gap for larger R is primarily due to conditioning and the fact that PG/NAG only project onto the equality constraint, whereas ADMM/CVX solve the full KKT system with box constraints.
+Intersection projections (e.g., Dykstra) could close this gap but would blur the simplicity advantage of PG/NAG; we leave such variants to future work
 Figure 3: Final-time relative $L^{2}$ error in the electric field versus stencil radius $R$ for each solver, using Crank–Nicolson time integration.
@@
 These results confirm that enforcing the skew-symmetry constraints during training yields learned stencils that are effectively energy-preserving under Crank–Nicolson time stepping.
+6.5 Dispersion analysis.
+To complement the end-time $L^{2}$ error metrics, we carry out a Fourier dispersion analysis of the semi-discrete Maxwell system with Crank–Nicolson time stepping.
+For a given stencil $w$ with Fourier symbol $\lambda(\theta)$, the one-step Crank–Nicolson update for each Fourier mode has eigenvalue
+$\mu(\theta)=\frac{1+\tfrac{\Delta t}{2}\,\lambda(\theta)}{1-\tfrac{\Delta t}{2}\,\lambda(\theta)}.$
+(a) Amplification error $|\mu(\theta)|-1$.
+(b) Phase dispersion $\arg\mu(\theta)/\theta$.
+Figure 5: Fourier dispersion analysis of the semi-discrete Maxwell system with Crank–Nicolson time stepping: (a) amplification error and (b) phase dispersion for learned vs.
+centered-difference stencils.
+We plot the amplification factor $|\mu(\theta)|$ and the normalized phase $\arg\mu(\theta)/\theta$ as functions of $\theta\in(0,\pi]$, and compare the learned energy-constrained stencils with standard centered-difference stencils of the same radius $R$; see Figs.
+5(a)5(b).
+Figure 5(a) shows that $|\mu(\theta)|\approx 1$ for all learned stencils and all wavenumbers, essentially indistinguishable from the centered-difference curves and the continuous reference.
+This confirms that the fully discrete schemes are non-dissipative and preserve the discrete electromagnetic energy in practice, consistent with the skew-adjointness constraints imposed on $D$.
+The phase plots in Fig.
+5(b) provide a more stringent test of accuracy.
+For $R=1,2,3$, the dispersion curves $\arg\mu(\theta)/\theta$ of the learned operators closely track those of the corresponding centered differences over the range of resolved wavenumbers, with only small deviations at the highest modes; in some mid-frequency bands the learned stencils are slightly closer to the continuous reference.
+Overall, this mode-by-mode dispersion analysis is more diagnostic than end-time $L^{2}$ errors and confirms that the hard skew-adjointness constraints produce learned stencils with physically meaningful wave-propagation properties.
+see, e.g., [21, 2]
+6.6 Learning a nonstandard skew-adjoint operator
+To mimic an effective spatial operator arising from unresolved heterogeneity or subgrid physics, we consider a “nonstandard” convolution stencil $w^{\star}\in\mathbb{R}^{2R+1}$ with radius $R=2$ that is skew-adjoint (i.e., satisfies the constraints (11)) but does not coincide with the usual fourth-order centered difference coefficients.
+In this setting the analytic form of the differential operator is not available, but we can still generate high-fidelity training trajectories by evolving Maxwell’s equations using $w^{\star}$.
+We then (i) learn an energy-constrained stencil $w_{\mathrm{QP}}$ by solving the quadratic program (20) with ADMM, and (ii) compare it to the standard centered-difference stencil $w_{\mathrm{CD}}$ of the same radius.
+Numerically we obtain
+$\frac{\|w_{\mathrm{QP}}-w^{\star}\|_{2}}{\|w^{\star}\|_{2}}\approx 6\times 10^{-8},\qquad\frac{\|w_{\mathrm{CD}}-w^{\star}\|_{2}}{\|w^{\star}\|_{2}}\approx 2.5\times 10^{-1},$
+so the learned operator recovers the nonstandard stencil to nearly machine precision, while the standard stencil incurs a $25\%$ relative coefficient error.
+Figure 6(a) compares the final-time electric field $E(x,T)$ produced by $w^{\star}$, $w_{\mathrm{CD}}$, and $w_{\mathrm{QP}}$.
+The centered-difference stencil exhibits a systematic phase and amplitude error, whereas the learned stencil is indistinguishable from the nonstandard operator.
+At the same time, all three stencils remain essentially energy-preserving under Crank–Nicolson time stepping (Figure 6(b)), demonstrating that the proposed method can capture effective, nonstandard differential operators while retaining a discrete Poynting theorem.
+(a) Noisy Discrete Energy Simulation
+(b) Noisy E(x,t)
+Figure 6: Nonstandard vs Standard Skew-adjoint operator
+6.7 Noisy training and ill-posed least squares
+$\min_{w\in\mathbb{R}^{2R+1}}\frac{1}{2}\|Aw-b\|_{2}^{2}+\frac{\lambda}{2}\|w\|_{2}^{2},$ (24)
+To emulate noisy or partially inconsistent experimental data, we add Gaussian noise to the time derivatives $dU/dt$ used in the regression problem (24).
+Specifically, we draw $\tilde{dU/dt}=dU/dt+\eta$ with $\eta$ i.i.d.
+$N(0,\sigma^{2})$ and form the corresponding least-squares system $(A,b_{\mathrm{noisy}})$.
+We then compute three stencils: (i) the standard centered difference $w_{\mathrm{CD}}$, (ii) an unconstrained least-squares stencil $w_{\mathrm{LS}}$ obtained by minimizing (24) with no constraints, and (iii) an energy-constrained stencil $w_{\mathrm{QP}}$ obtained by solving the quadratic program (20) with the same noisy data.
+Although $w_{\mathrm{LS}}$ achieves a very small training error, it violates the skew-adjointness constraints and develops very large high-frequency coefficients.
+When used in Crank–Nicolson time stepping, the resulting Maxwell semi-discretization is no longer energy-stable, and the discrete energy $\mathcal{E}^{n}$ blows up (Figure 7(a)).
+In contrast, the energy-constrained stencil $w_{\mathrm{QP}}$ remains exactly skew-adjoint by construction and yields stable dynamics with energy drift comparable to the hand-derived centered difference stencil.
+Figure 7(b) shows the final-time electric field $E(x,T)$ for all three operators.
+The solution obtained with $w_{\mathrm{QP}}$ is visually indistinguishable from the Central difference baseline, while the unconstrained LS stencil produces a completely corrupted field.
+This experiment illustrates a regime where training data are noisy or mildly inconsistent, and the energy constraints act as a form of physics-based regularization: they prevent overfitting to noise and guarantee a discrete Poynting theorem irrespective of the data quality.
+(a) Noisy Discrete Energy Simulation
+(b) Noisy E(x,t)
+Figure 7: Noisy training and ill-posed least squares
+Figure 8: Space-Time Plot of the Electric Field Under Noise Perturbation
 6.8 Solver convergence and runtime
@@
 8 Discussion and extensions
-The experiments above illustrate that simple linear constraints can encode important structural information in data-driven discretizations of PDEs.
-In this setting, the skew-symmetry constraint $Cw=d$ ensures semi-discrete energy conservation and leads to stencils that behave like standard centered differences but are obtained purely from data.
-Future work will learn staggered curl operators that preserve a discrete Poynting theorem in 2D/3D, integrate material models (Lorentz/Duffing) with structure-preserving polarization updates, and incorporate boundary treatments (ABC/CPML).
-On the optimization side, extending the comparison to higher-order stencils, non-quadratic objectives, and learned nonlinear fluxes would help clarify the regimes in which ADMM or interior-point methods offer the best tradeoff relative to accelerated first-order algorithms.
-Ultimately, combining learning with physics-motivated constraints promises accurate, stable reduced models for large-scale CEM while keeping the mathematical guarantees that make FDTD dependable [15, 5].
+Our results show that simple linear constraints can encode key physics in data-driven discretizations: enforcing $Cw=d$ yields skew-adjoint spatial operators that exactly preserve a semi-discrete energy identity, while remaining close to standard centered differences in the homogeneous 1D setting.
+To probe more realistic regimes, we also considered (i) a nonstandard target operator whose stencil does not match any classical centered-difference formula and showed that the constrained learner accurately recovers its space–time behavior, while the matching CD stencil incurs noticeable error, and (ii) a noisy regression experiment in which unconstrained least squares produces large coefficients and severe energy drift, whereas the energy- and box-constrained stencil stays well-conditioned, energy-stable, and faithful to the clean baseline $E(x,t)$.
+These examples suggest that physics-motivated constraints can regularize noisy or mildly mis-specified regression problems and make it feasible to learn effective Maxwell operators in settings where analytic stencils are difficult to derive or tune (e.g., heterogeneous media or band-limited/experimental data).
+A natural next step is to extend the same constrained-learning philosophy to staggered curl operators on Yee grids in 2D/3D, enforcing block skew-symmetry to preserve a discrete Poynting theorem, and to couple the learned operators with simple dispersive material models and absorbing layers in more realistic FDTD setups.
+Acknowledgements
+The author gratefully acknowledges Profs. Xiao Fu and Nathan Gibson for valuable guidance and feedback throughout this work.

### v4

@@
 We study data-driven construction of spatial discretizations for the one-dimensional Maxwell system.
-Using high-fidelity training data from a spectral discretization, we learn a linear convolution stencil that approximates the spatial derivative operator in Maxwell’s equations.
-We formulate a convex quadratic program for the stencil coefficients with linear constraints that enforce skew-adjointness of the discrete derivative; these constraints guarantee a semi-discrete electromagnetic energy identity and yield a CFL condition expressed directly in terms of the stencil’s Fourier symbol.
-We compare several convex solvers for the resulting quadratic program—projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point reference implemented in CVXPY—and evaluate the learned operators in time-dependent Maxwell simulations using a Crank–Nicolson (CN) discretization.
-Numerical experiments, including cases with nonstandard target operators and noisy training data, show that (i) energy-constrained learned stencils achieve accuracy comparable to standard central differences while exactly preserving the discrete electromagnetic energy under CN time-stepping, and (ii) ADMM and interior-point solvers produce nearly identical operators, with ADMM offering a favorable tradeoff between accuracy, constraint satisfaction, and runtime.
+Starting from high-fidelity spectral training data, we learn a linear convolution stencil that approximates the spatial derivative in Maxwell’s equations.
+The stencil coefficients are obtained from a convex quadratic program with linear constraints that enforce skew-adjointness of the discrete derivative; these constraints guarantee a semi-discrete electromagnetic energy identity and yield a CFL condition expressed directly in terms of the stencil’s Fourier symbol.
+We show that the learned operator has a purely imaginary symbol, induces purely oscillatory modal dynamics, and preserves the discrete Maxwell energy exactly under Crank–Nicolson time-stepping.
+Numerical experiments with smooth traveling-wave solutions demonstrate that the energy-constrained, ADMM-learned stencil with radius $R=1$ achieves approximately second-order spatial convergence in the PDE error, matching the expected order of the standard centered finite-difference stencil.
+We compare projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point solver for the resulting quadratic program, and we illustrate extensions to nonstandard target operators and noisy training data.
+Overall, the learned stencils achieve accuracy comparable to classical central differences while exactly preserving a discrete electromagnetic energy, making them viable drop-in, structure-preserving spatial operators for finite-difference time-domain methods and data-driven PDE solvers.
 1 Introduction
-Finite-difference time-domain (FDTD) methods are a standard tool for solving Maxwell’s equations in computational electromagnetics [21].
-Classical schemes, such as the Yee method, are derived by hand from the continuous equations, with spatial derivatives approximated using fixed central-difference stencils and stability enforced via analytical CFL conditions.
-Recently, there has been growing interest in data-driven discretizations, where discrete operators are learned from numerical or experimental data, for example via regression or neural networks.
-Examples include data-driven discretization of Burgers and Navier–Stokes equations and learned coarse-grid operators for PDEs [1].
-Most existing data-driven schemes aim to minimize approximation error but do not explicitly enforce structural properties such as stability, conservation, or energy balance.
-For Maxwell’s equations this is a significant limitation, since standard FDTD schemes are carefully designed so that discrete versions of divergence constraints and energy identities hold at the semi-discrete level.
+Finite-difference time-domain (FDTD) methods are a standard tool for solving Maxwell’s equations in computational electromagnetics [25, 26].
+Classical schemes, such as the Yee method [28], are derived directly from the continuous equations: spatial derivatives are approximated using fixed central-difference stencils, and stability is enforced via analytical CFL conditions.
+These schemes are carefully designed so that the discrete curl operator is skew-adjoint, which yields exact semi-discrete energy conservation and compatible discrete divergence constraints.
+Recently, there has been growing interest in data-driven discretizations, in which discrete operators are learned from numerical or experimental data, for example via regression or neural networks.
+Examples include data-driven finite-difference discretizations for Burgers’ equation and the Navier–Stokes equations [1], data-driven closures for turbulent flows [21, 14], and sparse-regression or neural approaches for discovering governing equations and hidden dynamics [24, 22, 7].
+These methods typically aim to minimize approximation error on a chosen training distribution, but they do not, in general, explicitly enforce structural properties such as stability, conservation, or energy balance.
+For Maxwell’s equations, this is a significant limitation.
+Classical FDTD schemes are constructed so that discrete versions of divergence constraints and energy identities hold at the semi-discrete level, ensuring that numerical solutions respect key physical invariants.
+Learned operators that ignore these structures risk exhibiting uncontrolled numerical dissipation or growth, even if they fit the training data well.
 In this work we address this gap in a simple but concrete setting: the one-dimensional Maxwell system with periodic boundary conditions. Our goal is to learn a discrete spatial derivative operator
-$D:\mathbb{R}^{N}\to\mathbb{R}^{N}$
+$D:\mathbb{R}^{N}\to\mathbb{R}^{N},$
 represented as a convolution stencil, directly from data.
 We impose linear constraints on the stencil that enforce skew-adjointness of $D$ with respect to the discrete $L^{2}$ inner product.
-This structural constraint guarantees exact conservation of discrete electromagnetic energy at the semi-discrete level, regardless of the details of the training data.
-Classical FDTD schemes such as the Yee method [21] are hand-derived so that the discrete curl operator is skew-adjoint, yielding exact semi-discrete energy conservation.
-High-order extensions and operator-splitting schemes similarly maintain this structure [3, 6].
-Data-driven PDE schemes learn local stencils or coarse-grid operators directly from numerical data.
-Bar-Sinai et al.
-learn finite-difference discretizations for Burgers and Navier–Stokes, showing that learned operators can match or exceed classical schemes on specific training distributions [1].
-Many follow-up works learn PDE operators or closure models [17, 11, 20, 18, 9], but typically do not enforce stability, skew-adjointness, or conservation laws.
-Recent neural-operator surrogates for electromagnetics incorporate physics through architectural biases or penalty terms.
-Dynamic-convolution neural operators for FDTD-like EM propagation impose causal receptive fields and material dispersion structure [14].
+This structural constraint guarantees exact conservation of discrete electromagnetic energy at the semi-discrete level, independently of the details of the training data or the optimization procedure.
+Our approach is inspired by two lines of work.
+On the one hand, classical structure-preserving discretizations for Maxwell and related PDEs use staggered grids and skew-adjoint discrete curl operators to obtain discrete energy identities and long-time stability [4, 8, 19].
+High-order and operator-splitting variants maintain these properties by design [4, 3].
+On the other hand, data-driven PDE schemes learn local stencils or coarse-grid operators from high-fidelity simulations, often achieving accuracy comparable to or better than hand-crafted schemes on specific tasks [1, 11, 6].
+However, these learned operators usually encode physical structure only implicitly, through the choice of loss function, training data, or architecture, rather than via hard constraints.
+Related ideas also appear in recent neural-operator surrogates for electromagnetics and other time-dependent PDEs.
+Dynamic-convolution neural operators for FDTD-like EM propagation impose causal receptive fields and material-dispersion structure [17, 13].
 Domain-decomposition neural operators couple learned subdomain solvers through interface penalties or consistency conditions [18, 10].
-Other neural operator frameworks embed conservation via soft constraints or corrective layers [13, 12].
-These approaches improve stability but enforce physical structure only approximately, through optimization.
+Other frameworks embed conservation laws via soft constraints or corrective layers [16, 15, 27].
+These approaches can significantly improve stability and physical fidelity, but they generally enforce conservation or energy balance only approximately, depending on optimization and hyperparameters.
 Building on this context, our main contributions are as follows.
-First, we formulate a convex quadratic optimization problem for learning convolution stencils for the 1D Maxwell system from spectral training data, with an objective given by a least-squares misfit between the learned discrete derivative and high-fidelity derivatives, and linear constraints that encode skew-symmetry of the stencil coefficients together with simple box bounds.
-Second, we show that these linear constraints are exactly equivalent to skew-adjointness of the corresponding convolution operator with respect to the discrete $L^{2}$ inner product, and hence to a semi-discrete electromagnetic energy identity, yielding what is, to our knowledge, the first learned Maxwell stencil with a guaranteed semi-discrete energy conservation law.
-Third, we characterize the numerical wave speed and associated CFL restriction of the learned discrete curl operator purely in terms of its Fourier symbol $\mu(\theta)$.
-Finally, we implement projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point method (via CVXPY) to solve the constrained quadratic program, and we benchmark the resulting stencils in time-dependent 1D Maxwell simulations; the experiments compare PDE error, energy-constraint residuals, and energy evolution across stencil radii, and demonstrate that ADMM and interior-point methods produce nearly identical, highly accurate stencils, with ADMM offering a favorable tradeoff between accuracy, constraint satisfaction, and runtime.
-Although our experiments focus on a simple 1D periodic model, the framework and energy constraints extend naturally to higher-dimensional settings and more complex material models.
+We formulate a convex quadratic optimization problem for learning convolution stencils for the 1D Maxwell system from spectral training data.
+The objective is a least-squares misfit between the learned discrete derivative and high-fidelity derivatives, subject to linear constraints that encode skew-adjointness of the stencil coefficients together with simple box bounds.
+We show that these linear constraints are exactly equivalent to skew-adjointness of the corresponding convolution operator with respect to the discrete $L^{2}$ inner product, which in turn yields an exact semi-discrete electromagnetic energy identity and expresses the CFL condition directly in terms of the stencil’s Fourier symbol $\mu(\theta)$.
+This yields, to our knowledge, the first learned Maxwell stencil with a guaranteed semi-discrete energy conservation law.
+We characterize the numerical wave speed and associated CFL restriction of the learned discrete curl operator via this symbol.
+We implement projected gradient, Nesterov-accelerated gradient, ADMM, and an interior-point method to solve the constrained quadratic program, and we benchmark the resulting stencils in time-dependent 1D Maxwell simulations.
+We also illustrate how the same constrained learning framework can recover nonstandard, coarse-grid effective operators and regularize synthetic spectral training data contaminated by noise or mild inconsistencies.
+In particular, the ADMM-learned radius-$R=1$ stencil exhibits approximately second-order spatial convergence for smooth traveling-wave solutions, matching the expected order of the standard centered finite-difference stencil it approximates.
+The experiments compare PDE error, energy-constraint residuals, and energy evolution across stencil radii, and demonstrate that ADMM and the interior-point solver produce nearly identical, highly accurate stencils, with ADMM offering a favorable tradeoff between accuracy, constraint satisfaction, and runtime.
 Our results illustrate how even simple linear constraints in a data-driven discretization can encode important physics and restore structural properties such as energy conservation.
-Our method imposes hard linear constraints that make the learned convolution stencil exactly skew-adjoint, guaranteeing a discrete electromagnetic energy identity independent of training accuracy or solver choice.
-This places our work closest to Bar-Sinai et al.
-but with strict structure preservation rather than penalty-based regularization, and complementary to physics-informed neural operators, which typically enforce conservation approximately.
-To the best of our knowledge, this is the first work to combine data-driven stencil learning for Maxwell’s equations with linear constraints that (i) exactly characterize skew-symmetric convolution operators and (ii) yield a rigorous semi-discrete energy identity and a CFL condition expressed directly in terms of the learned stencil’s Fourier symbol.
-This provides a bridge between unconstrained data-driven discretizations and classical energy-preserving FDTD schemes, while remaining compatible with standard convex optimization toolchains (projected gradient, Nesterov acceleration, ADMM, and interior-point solvers).
-Beyond the homogeneous model, Sections 6.6 and 6.7 illustrate how the same constrained learning framework can recover nonstandard, effective operators and regularize noisy or mildly inconsistent training data.
+To the best of our knowledge, this is the first data-driven convolution stencil for the one-dimensional Maxwell system whose learned coefficients are constrained to guarantee a semi-discrete electromagnetic energy identity.
+The paper is organized as follows.
+Sections 2–3 introduce the semi-discrete 1D Maxwell system, derive the corresponding discrete energy identity, and analyze the numerical wave speed and CFL restriction associated with the learned operator.
+Sections 4–6 describe the learning formulation and optimization algorithms, and present numerical experiments that evaluate accuracy, energy behavior, dispersion, and robustness to nonstandard targets and noisy training data.
 2 Semi-discrete Maxwell system and energy identity
@@
 $(D\mathbf{u})_{i}=\sum_{\ell=-R}^{R}w_{\ell}\,u_{i+\ell}=\sum_{j=0}^{N-1}w_{i-j}\,u_{j},$
-where $j=i+\ell$ (modulo NN). Hence $D_{ij}=w_{i-j}$.[3]
+where $j=i+\ell$ (modulo NN). Hence $D_{ij}=w_{i-j}$; see, e.g., [4]. ∎
 Theorem 2 (Characterization of skew-symmetric stencils).
-[3]Let $D$ be the convolution operator (2) with periodic boundary conditions and stencil $w\in\mathbb{R}^{2R+1}$.
+Let $D$ be the convolution operator (2) with periodic boundary conditions and stencil $w\in\mathbb{R}^{2R+1}$.
 Then $D$ is skew-adjoint with respect to (4), i.e., $D^{\top}=-D$, if and only if the stencil satisfies
@@
 which expresses the admissible time step solely in terms of the learned symbol (12).
-In this way the learned stencil can be interpreted within the familiar CFL framework for explicit FDTD schemes, although in our numerical experiments we will use an implicit time integrator.In this work we focus on CN, which is unconditionally stable for the skew-adjoint semi-discrete operator.
-A systematic comparison with leapfrog using the derived CFL bound is left for future work
+In this way the learned stencil can be interpreted within the familiar CFL framework for explicit FDTD schemes, although in our numerical experiments we will use an implicit time integrator.
+In the present work we focus on CN, which is unconditionally stable for the skew-adjoint semi-discrete operator.
+A systematic comparison with leapfrog using the derived CFL bound is left for future work.
 3.1 Crank–Nicolson time integration
@@
 5.2 ADMM
-To better handle the combination of equality and box constraints, we also implement an alternating direction method of multipliers (ADMM) [4], [5] scheme.
+To better handle the combination of equality and box constraints, we also implement an alternating direction method of multipliers (ADMM) [5] scheme.
 We introduce a copy $z$ of $w$ and enforce the box constraint on $z$ only:
@@
 We again monitor the objective, equality residual, and primal difference $\|w^{k+1}-w^{k}\|$.
-5.3 Interior-point reference (CVXPY)
+5.3 Interior-point reference
 Finally, we use the CVXPY modeling framework with a cone solver (SCS in our experiments) to obtain a reference solution to (19).
@@
 The ADMM and projected-gradient/Nesterov solvers are written explicitly as described in Section 5.
-The interior-point reference solution is computed using CVXPY (version X.Y) with the SCS solver.
-Runtimes are reported for a single-threaded run on a (Apple M4, 16 GB RAM) using Python 3.11.
+The interior-point reference solution is computed using CVXPY with the SCS solver.
+Runtimes are reported for a single-threaded run on an Apple M4 laptop with 16 GB RAM using Python 3.11.
 Although absolute times depend on hardware and implementation details, the relative trends between first-order methods (PG/NAG) and the second-order ADMM/CVX approach remained consistent across several runs.
@@
 and we compute a reference solution using the exact central-difference stencil
-$w_{\mathrm{exact}}=\frac{1}{2\Delta x}(-1,\,0,\,1),$ (22)
+$w_{\mathrm{exact}}=\frac{1}{2\Delta x}(-1,\,0,\,1).$ (22)
 We then integrate the Maxwell system with each learned stencil using CN and measure the relative $L^{2}$ error in the electric field at final time:
@@
 over time. For the exact stencil and CN time-stepping, the code confirms that $\mathcal{E}^{n}$ is constant to within floating-point roundoff, consistent with the analysis in Section 2.
-6.2 Learned stencils for $R=1$
+6.2 Learned stencils
 Table 1 shows the learned stencils obtained by each solver compared with $w_{\mathrm{exact}}$.
@@
 CVX 32.906115-32.906115 $1.57\!\times\!10^{-14}$ $32.906115$
-Table 1: Learned stencils for radius $R=1$.
-In In Figures 1 and 2, we compare the space-time evolution and the final-time profile of the electric field $E(x,t)$ for each learned stencil against the reference FD solution for R =1.
+Table 1: Learned stencils for radius $R=1$ compared with the exact central-difference stencil.
+In Figures 1 and 2, we compare the space-time evolution and the final-time profile of the electric field $E(x,t)$ for each learned stencil against the reference FD solution.
 The final-time and space-time profiles in Figure 2 are visually indistinguishable from the reference solution.
-Figure 1: Space-Time Evolution
+Figure 1: Space-time evolution of the electric field $E(x,t)$ for the learned stencils and the exact central-difference stencil.
 Figure 2: Final-time electric field $E(x,T)$ for the exact central-difference stencil and the learned stencils for $R=1$ under Crank–Nicolson time-stepping.
@@
 This $R=1$ case therefore provides a baseline for the higher-radius stencils and noisy-training experiments considered below.
-6.3 Error and energy residual versus stencil radius
-To explore the effect of stencil size, we repeat the learning and time-integration procedure for radii $R=1,2,3,4$.
-Figure 3 shows the final-time relative $L^{2}$ error $\mathrm{err}(w)$ for each method as a function of $R$.
-We initialize all optimizations at the zero stencil; even from this uninformative start, the energy-constrained learner recovers stencils close to classical centered differences.
-Figure 3 shows that for $R=1$ all four solvers achieve comparable accuracy, with errors on the order of $5\times 10^{-3}$.
-As the radius increases, the behavior diverges: the projected-gradient and Nesterov methods exhibit substantially larger errors, with PG reaching $\mathcal{O}(1)$ error and NAG reaching $\mathcal{O}(10^{-2})$ for $R=3$.
-In contrast, the ADMM and CVX solutions maintain errors in a narrow band between approximately $3\times 10^{-3}$ and $6\times 10^{-3}$ for all radii tested.
-This indicates that the second-order KKT solve in ADMM and the interior-point method in CVX are more effective at exploiting the additional degrees of freedom introduced by larger stencils, whereas the first-order projected methods struggle to fully optimize the objective within the available iteration budget.We believe the gap for larger R is primarily due to conditioning and the fact that PG/NAG only project onto the equality constraint, whereas ADMM/CVX solve the full KKT system with box constraints.
-Intersection projections (e.g., Dykstra) could close this gap but would blur the simplicity advantage of PG/NAG; we leave such variants to future work
-Figure 3: Final-time relative $L^{2}$ error in the electric field versus stencil radius $R$ for each solver, using Crank–Nicolson time integration.
+6.3 Spatial convergence for the ADMM–learned stencil
+We consider the 1D Maxwell system in normalized units,
+$\partial_{t}E=\partial_{x}H,\qquad\partial_{t}H=\partial_{x}E,$ (24)
+with $\varepsilon_{0}=\mu_{0}=1$ so that the wave speed is $c=1$. On the periodic domain $x\in[0,L]$ we choose the traveling-wave exact solution
+$E_{\mathrm{ex}}(x,t)=\sin\!\left(\frac{2\pi}{L}(x-t)\right),\qquad H_{\mathrm{ex}}(x,t)=\cos\!\left(\frac{2\pi}{L}(x-t)\right),$ (25)
+which satisfies (24) and is $L$-periodic in space.
+The initial data at $t=0$ are taken from (25), and the numerical solution is advanced to a final time $T=10$ using the Crank–Nicolson time integrator with a spatial derivative given by the ADMM–learned stencil of radius $R=1$.
+Let $E^{N}_{i}$ denote the numerical approximation to $E_{\mathrm{ex}}(x_{i},T)$ at the grid points $x_{i}=i\,\Delta x$, $i=0,\dots,N_{x}-1$.
+We measure the relative discrete $L^{2}$ error in the electric field by
+$\mathrm{err}_{E}(\Delta x)=\frac{\left\|E^{N}-E_{\mathrm{ex}}(\cdot,T)\right\|_{\ell^{2}}}{\left\|E_{\mathrm{ex}}(\cdot,T)\right\|_{\ell^{2}}},\qquad\|v\|_{\ell^{2}}^{2}=\Delta x\sum_{i=0}^{N_{x}-1}v_{i}^{2}.$ (26)
+The timestep is chosen proportionally to the mesh width, $\Delta t=0.2\,\Delta x$, so that we probe the spatial convergence of the learned stencil.
+Table 2: Convergence of the ADMM–learned stencil with radius $R=1$. Relative $L^{2}$ error in $E(\cdot,T)$
+$N_{x}$ $\Delta x$ rel. $L^{2}(E)$ error order
+64 $1.562\times 10^{-2}$ $2.097\times 10^{0}$
+128 $7.812\times 10^{-3}$ $4.875\times 10^{-1}$ $2.10$
+256 $3.906\times 10^{-3}$ $1.192\times 10^{-1}$ $2.03$
+512 $1.953\times 10^{-3}$ $585\times 10^{-2} 2.2$02.20
+The table2 confirms approximately second-order spatial accuracy for the ADMM–learned radius-$1$ stencil.
+Figure 3: Log–log plot of the relative $L^{2}$ error $\mathrm{err}_{E}(\Delta x)$ in $E(\cdot,T)$ versus $\Delta x$ for the ADMM–learned stencil.
 6.4 Discrete energy conservation for learned Maxwell stencils
@@
 These results confirm that enforcing the skew-symmetry constraints during training yields learned stencils that are effectively energy-preserving under Crank–Nicolson time stepping.
-6.5 Dispersion analysis.
+6.5 Dispersion analysis
 To complement the end-time $L^{2}$ error metrics, we carry out a Fourier dispersion analysis of the semi-discrete Maxwell system with Crank–Nicolson time stepping.
@@
 For $R=1,2,3$, the dispersion curves $\arg\mu(\theta)/\theta$ of the learned operators closely track those of the corresponding centered differences over the range of resolved wavenumbers, with only small deviations at the highest modes; in some mid-frequency bands the learned stencils are slightly closer to the continuous reference.
-Overall, this mode-by-mode dispersion analysis is more diagnostic than end-time $L^{2}$ errors and confirms that the hard skew-adjointness constraints produce learned stencils with physically meaningful wave-propagation properties.
-see, e.g., [21, 2]
+Overall, this mode-by-mode dispersion analysis is more diagnostic than end-time $L^{2}$ errors and confirms that the hard skew-adjointness constraints produce learned stencils with physically meaningful wave-propagation properties; see, e.g., [25, 2].
 6.6 Learning a nonstandard skew-adjoint operator
@@
 At the same time, all three stencils remain essentially energy-preserving under Crank–Nicolson time stepping (Figure 6(b)), demonstrating that the proposed method can capture effective, nonstandard differential operators while retaining a discrete Poynting theorem.
-(a) Noisy Discrete Energy Simulation
-(b) Noisy E(x,t)
-Figure 6: Nonstandard vs Standard Skew-adjoint operator
+(a) Final-time electric field $E(x,T)$ for nonstandard and learned stencils.
+(b) Discrete energy error $\mathcal{E}^{n}-\mathcal{E}^{0}$.
+Figure 6: Nonstandard vs. standard skew-adjoint operators: learned stencil $w_{\mathrm{QP}}$ compared with $w^{\star}$ and the centered difference $w_{\mathrm{CD}}$.
 6.7 Noisy training and ill-posed least squares
-$\min_{w\in\mathbb{R}^{2R+1}}\frac{1}{2}\|Aw-b\|_{2}^{2}+\frac{\lambda}{2}\|w\|_{2}^{2},$ (24)
 To emulate noisy or partially inconsistent experimental data, we add Gaussian noise to the time derivatives $dU/dt$ used in the regression problem (17).
-Specifically, we draw $\tilde{dU/dt}=dU/dt+\eta$ with $\eta$ i.i.d.
+Specifically, we draw $\widetilde{dU/dt}=dU/dt+\eta$ with $\eta$ i.i.d.
 $N(0,\sigma^{2})$ and form the corresponding least-squares system $(A,b_{\mathrm{noisy}})$.
@@
 Figure 7(b) shows the final-time electric field $E(x,T)$ for all three operators.
-The solution obtained with $w_{\mathrm{QP}}$ is visually indistinguishable from the Central difference baseline, while the unconstrained LS stencil produces a completely corrupted field.
+The solution obtained with $w_{\mathrm{QP}}$ is visually indistinguishable from the central-difference baseline, while the unconstrained LS stencil produces a completely corrupted field.
 This experiment illustrates a regime where training data are noisy or mildly inconsistent, and the energy constraints act as a form of physics-based regularization: they prevent overfitting to noise and guarantee a discrete Poynting theorem irrespective of the data quality.
-(a) Noisy Discrete Energy Simulation
-(b) Noisy E(x,t)
-Figure 7: Noisy training and ill-posed least squares
-Figure 8: Space-Time Plot of the Electric Field Under Noise Perturbation
+(a) Discrete energy error $\mathcal{E}^{n}-\mathcal{E}^{0}$ under noisy training.
+(b) Final-time electric field $E(x,T)$ for noisy training.
+Figure 7: Noisy training and ill-posed least squares: comparison of unconstrained LS, central difference, and energy-constrained stencils.
+Figure 8: Space-time plot of the electric field $E(x,t)$ under noisy training perturbations.
 6.8 Solver convergence and runtime
 Finally, we compare the convergence histories of the different solvers.
-Figure 9(a) plots the objective value versus iteration for each first-order method for a fixed radius.
-Figure 9(b) shows the same objective versus wall-clock time.
+Figure 9 plots the objective value versus iteration and wall-clock time for each first-order method for a fixed radius.
 These plots illustrate the tradeoffs among PG, NAG, and ADMM in terms of iterations, per-iteration cost, and constraint enforcement.
-(a) Objective vs iteration
-(b) Objective vs runtime
+(a) Objective vs. iteration.
+(b) Objective vs. runtime.
 Figure 9: Convergence behavior of PG, NAG, and ADMM for a fixed stencil radius $R$. Left: objective value versus iteration count. Right: objective value versus wall-clock runtime.
@@
 Our results show that simple linear constraints can encode key physics in data-driven discretizations: enforcing $Cw=d$ yields skew-adjoint spatial operators that exactly preserve a semi-discrete energy identity, while remaining close to standard centered differences in the homogeneous 1D setting.
-To probe more realistic regimes, we also considered (i) a nonstandard target operator whose stencil does not match any classical centered-difference formula and showed that the constrained learner accurately recovers its space–time behavior, while the matching CD stencil incurs noticeable error, and (ii) a noisy regression experiment in which unconstrained least squares produces large coefficients and severe energy drift, whereas the energy- and box-constrained stencil stays well-conditioned, energy-stable, and faithful to the clean baseline $E(x,t)$.
+To probe more realistic regimes, we also considered a nonstandard target operator whose stencil does not match any classical centered-difference formula and showed that the constrained learner accurately recovers its space–time behavior, while the matching CD stencil incurs noticeable error, and a noisy regression experiment in which unconstrained least squares produces large coefficients and severe energy drift, whereas the energy- and box-constrained stencil stays well-conditioned, energy-stable, and faithful to the clean baseline $E(x,t)$.
 These examples suggest that physics-motivated constraints can regularize noisy or mildly mis-specified regression problems and make it feasible to learn effective Maxwell operators in settings where analytic stencils are difficult to derive or tune (e.g., heterogeneous media or band-limited/experimental data).
-A natural next step is to extend the same constrained-learning philosophy to staggered curl operators on Yee grids in 2D/3D, enforcing block skew-symmetry to preserve a discrete Poynting theorem, and to couple the learned operators with simple dispersive material models and absorbing layers in more realistic FDTD setups.
+A natural next step is to extend the same constrained-learning philosophy to staggered curl operators on Yee grids in higher dimensions, enforcing block skew-symmetry to preserve a discrete Poynting theorem, and to couple the learned operators with simple dispersive material models and absorbing layers in more realistic FDTD setups.
 Acknowledgements
 The author gratefully acknowledges Profs. Xiao Fu and Nathan Gibson for valuable guidance and feedback throughout this work.
+Appendix: Spectral structure of the learned stencil
+Theorem 3 (Spectral structure and modal energy conservation for the learned stencil).
+Let $w^{\star}\in\mathbb{R}^{2R+1}$ be any solution of the quadratic program (19) (or the equality-constrained variant (20)), and let $D^{\star}$ be the associated periodic convolution operator defined by (2).
+Let $\mu^{\star}(\theta)$ denote the Fourier symbol of $D^{\star}$,
+$\mu^{\star}(\theta):=\sum_{k=-R}^{R}w_{k}^{\star}\,\mathrm{e}^{\mathrm{i}k\theta},\qquad\theta\in[-\pi,\pi].$
+Then the following hold:
+  1. (i)
+    Because $w^{\star}$ satisfies the linear constraint $Cw^{\star}=d$, it is skew-symmetric in the sense of (10), i.e.
+    $w_{0}^{\star}=0,\qquad w_{-\ell}^{\star}=-w_{+\ell}^{\star},\quad\ell=1,\dots,R.$
+    Consequently, the symbol $\mu^{\star}(\theta)$ is purely imaginary and odd:
+    $\mu^{\star}(\theta)\in\mathrm{i}\mathbb{R},\qquad\mu^{\star}(-\theta)=-\mu^{\star}(\theta)\quad\text{for all }\theta\in[-\pi,\pi].$
+  2. (ii)
+    Consider the semi-discrete Maxwell system with the learned operator $D^{\star}$,
+    $\frac{d}{dt}\begin{bmatrix}\mathbf{H}\\[2.5pt] \mathbf{E}\end{bmatrix}=A^{\star}\begin{bmatrix}\mathbf{H}\\[2.5pt] \mathbf{E}\end{bmatrix},\qquad A^{\star}=\begin{bmatrix}0&D^{\star}\\ (D^{\star})^{\top}&0\end{bmatrix}$,
+    together with the discrete $L^{2}$ inner product (4) and energy (5).
+Then $A^{\star}$ is skew-adjoint and diagonalizable in the discrete Fourier basis.
+For each discrete wavenumber $\theta$, the corresponding $2\times 2$ block has eigenvalues
+    $\lambda_{\pm}^{\star}(\theta)=\pm\,\mathrm{i}\,|\mu^{\star}(\theta)|.$
+    In particular, each Fourier mode evolves as a harmonic oscillator with frequency $|\mu^{\star}(\theta)|$.
+  3. (iii)
+    The discrete electromagnetic energy associated with the learned stencil can be written as a sum of modal contributions and is conserved mode by mode.
+If $\widehat{E}^{\star}(\theta,t)$ and $\widehat{H}^{\star}(\theta,t)$ denote the discrete Fourier coefficients of $\mathbf{E}(t)$ and $\mathbf{H}(t)$, then
+    $\mathcal{E}(t)=\frac{\Delta x}{2}\sum_{\theta}\bigl(|\widehat{E}^{\star}(\theta,t)|^{2}+|\widehat{H}^{\star}(\theta,t)|^{2}\bigr),$
+    and for each $\theta$ one has
+    $\frac{d}{dt}\bigl(|\widehat{E}^{\star}(\theta,t)|^{2}+|\widehat{H}^{\star}(\theta,t)|^{2}\bigr)=0.$
+    Thus the total discrete energy and each modal energy contribution are preserved exactly by the semi-discrete learned Maxwell system.
+Proof.
+Because $w^{\star}$ is a solution of (19) or (20), it satisfies the linear constraint $Cw^{\star}=d$ by construction.
+By Corollary 1, this is equivalent to the skew-symmetry conditions (10), which immediately imply the structure of $\mu^{\star}(\theta)$ and the identities in part (i) via the same calculation as in Theorem 2.
+Skew-symmetry of $w^{\star}$ implies $(D^{\star})^{\top}=-D^{\star}$, so the block matrix $A^{\star}$ is skew-adjoint with respect to the discrete inner product (4).
+Since $D^{\star}$ is a periodic convolution, it is diagonalized by the discrete Fourier transform; the $2\times 2$ block for each mode $\theta$ has the form
+$A^{\star}(\theta)=\begin{bmatrix}0&\mu^{\star}(\theta)\\ -\mu^{\star}(\theta)&0\end{bmatrix}$,
+whose eigenvalues are $\lambda_{\pm}^{\star}(\theta)=\pm\,\mathrm{i}\,|\mu^{\star}(\theta)|$, yielding part (ii).
+Finally, unitarity of the discrete Fourier transform implies that the discrete $L^{2}$ norms of $\mathbf{E}$ and $\mathbf{H}$ are equal to the sums of squared modal amplitudes, so the energy $\mathcal{E}(t)$ decomposes as in part (iii).
+For each fixed $\theta$, the $2\times 2$ system governed by the skew-Hermitian matrix $A^{\star}(\theta)$ preserves the Euclidean norm of $(\widehat{H}^{\star}(\theta,t),\widehat{E}^{\star}(\theta,t))$, which implies modal energy conservation and hence conservation of the total discrete energy.
+∎